### v2

# QUIDS: Query Intent Generation via Dual Space Modeling

Abstract.
Query understanding is a crucial component of Information Retrieval (IR), aimed at identifying the underlying search intent of textual queries. However, most existing approaches oversimplify this task into query classification or clustering, which fails to fully capture the nuanced intent behind the query. In this paper, we address the task of query intent generation: to automatically generate detailed and precise intent descriptions for search queries using relevant and irrelevant documents given a query. These intent descriptions can help users understand why the search engine considered the top-ranked documents relevant, and provide more transparency to the retrieval process. We propose a dual-space model that uses semantic relevance and irrelevance information in the returned documents to explain the understanding of the query intent. Specifically, in the encoding process, we project, separate, and distinguish relevant and irrelevant documents in the representation space. Then, we introduce a semantic decoupling model in the novel disentangling space, where the semantics of irrelevant information are removed from the relevant space, ensuring that only the essential and relevant intent is captured. This process refines the understanding of the query and provides more accurate explanations for the search results. Experiments on benchmark data demonstrate that our methods produce high-quality query intent descriptions, outperforming existing methods for this task, as well as state-of-the-art query-based summarization methods. A token-level visualization of attention scores reveals that our model effectively reduces the focus on irrelevant intent topics. Our findings open up promising research and application directions for query intent generation, particularly in exploratory search.
Information Retrieval, Query intent generation, Contrastive learning, Transformers
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Users and interactive retrievalccs: Information systems Query intentccs: Computing methodologies Natural language generation
1. Introduction
Query understanding is a focal point in the Information Retrieval (IR) community, aimed at determining the underlying search intent of a textual query input. This facet of IR is crucial for several reasons. Firstly, it enables search engines to distinguish and filter out unwanted interpretations of the current search query, ensuring relevance and accuracy in results. This is important because queries tend to be short and underspecified and are therefore often ambiguous (Guo and Lan, 2020). Secondly, users benefit when search engines provide feedback, such as query expansions or more descriptive query intents (Chen et al., 2021). This enhances the search experience by enabling users to refine and adjust their queries based on insights derived from the provided query understanding.
Figure 1. An application of the query intent generation task as a form of explanation to the user in exploratory search. The top-ranked documents are used as pseudo-relevant documents to generate the intent. Information relevant to forming the intent is highlighted in boldface.
Existing methods for query understanding can be broadly categorized into three groups. Query classification methods (Broder, 2002; Verberne et al., 2013) aim to categorize queries into distinct classes based on intent taxonomies. However, many of these methods tend to abstract intents into broader categories, lacking granularity in their analysis. Second, query clustering methods (Wen et al., 2002; Hong et al., 2016) aim to group query intents into clusters, yet encounter challenges in precisely interpreting individual queries within these clusters. The third approach, which is the most effective way to capture query intent, is query intent description generation, where the intent is described in textual format. However, this task has been largely overlooked in existing work, with the most recent contribution by (Zhang et al., 2020), which still relies on RNN-based methods and does not leverage the advancements in transformer models.
Figure 1 illustrates an example of query intent generation for the query ‘Freon-12’, which has an exploratory intent – aimed at exploring topics without a clearly defined goal  (Palagi et al., 2017). Since the true intent behind the query is unknown to the search engine, it relies on the retrieved documents to infer it. Thus, it treats the top-ranked documents as pseudo-relevant and the low-ranked documents as irrelevant. The intent description generated from both sets of documents is highlighted with a green background, while the information directly contributing to intent formation is marked in boldface. From the relevant documents, key terms like ‘Freon-12’, ‘refrigerant’, and ‘phased out’ are captured and emphasized in the intent description. In contrast, topics such as ‘costs’ and ‘prices’, which appear in both relevant and irrelevant documents, are deprioritized and excluded from the final intent description. The resulting intent generation illustrates how the search engine interprets the user’s query based on the retrieved documents. Furthermore, this approach allows users to reflect on their interests more clearly, potentially guiding them to reformulate their query more precisely and achieve more targeted search results. This iterative process of query refinement helps bridge the gap between exploratory search and more specific information needs, enhancing the overall search experience.
In scenarios where the ground truth intent is available, a related task to train models for generating query intent is query-focused summarization (QFS). This task is to generate summaries tailored to specific user queries or information needs, drawing from pseudo-relevant (high-ranked) documents. However, the focus in QFS is on summarizing content rather than explicitly identifying or reproducing query intents. Typically, QFS does not involve irrelevant information, which means the resulting summaries may include unfiltered or irrelevant details that are not truly reflective of the query’s intent. On the other hand, previous work (Zhang et al., 2020) on the Q2ID task incorporates irrelevant documents for contrast, but it does not establish a direct connection between the query and irrelevant documents. Our approach explicitly models both relevant and irrelevant intent spaces in a query-aware manner. Specifically, we construct the final intent space by subtracting the irrelevant intent space from the relevant one. This subtraction ensures that the final intent description is closely aligned with the query while effectively filtering out irrelevant information.
We propose a novel pipeline, implementing contrastive learning through dual space modeling in transformer-based (Vaswani et al., 2017) models, particularly T5 (Raffel et al., 2020a) and BART (Lewis et al., 2020). Specifically, we use a dual-encoder architecture and perform contrastive learning in the representation space using encoder embeddings of both relevant and irrelevant documents. During the decoding phase, we adapt the transformer decoder to conduct contrastive learning using decoder hidden states in the newly proposed disentangling space, where we model query-aware relevant and irrelevant intent spaces. This dual-space contrastive learning approach ensures that the model learns to differentiate between relevant and irrelevant documents, while also guiding it to produce high-quality intent descriptions with attention to the learned relevant and irrelevant intent space. Moreover, we introduce a method to augment hard negative documents driven by intent, to expand the irrelevant information representation space during training.
We evaluate our model using the benchmark dataset Q2ID, sourced from TREC and SemEval, which includes two types of query intent. Our results demonstrate that our model outperforms state-of-the-art baselines in terms of ROUGE metrics and BERTScore. Additionally, we perform human evaluations and LLM-based evaluations using tailored metrics, and provide a comprehensive analysis of our approach. To further illustrate the effect of irrelevant documents, we conduct a case study in which we highlight the information used to generate a semantically rich intent token based on the decoder’s cross-attention weights at the token level for both relevant and irrelevant documents. In summary, our contributions are as follows:
  • We introduce contrastive learning in both the representation space and the disentangling space of transformer models, effectively capturing contrasting information from relevant and irrelevant documents.
  • Our model generates high-quality intent descriptions, with performance significantly enhanced by incorporating hard negative data augmentation during training.
  • We perform a thorough evaluation of our model, providing us with insights into the model’s strengths and weaknesses, as well as its potential application scenarios.
2. Related Work
2.1. Query Understanding
The most closely related tasks in query understanding include query classification, query clustering, and query expansion. Query classification (Guo and Lan, 2020; Li et al., 2023) aims to classify queries into one or more pre-defined target taxonomies, based on factors like intent (Broder, 2002; Rose and Levinson, 2004), topic (Li et al., 2005; Beitzel et al., 2005), information type (Pasca and Harabagiu, 2001) or other classification dimensions. However, due to the typically short and ambiguous nature of search queries, as well as their evolving meanings, these taxonomies may lack sufficient granularity and become outdated over time (Guo and Lan, 2020). This can lead to the loss of detailed information in the classification process. Query clustering attempts to discover search topics or intents by mining clusters of queries. The similarity between queries is typically measured using data from user click-through logs (Wen et al., 2002; Rani and Babu, 2019; Lu et al., 2024) or top-ranked documents (Hong et al., 2016). In a recent application, query clustering has been employed in search engines to display real-time trending user queries, capturing shifts in queries and news articles over consecutive timestamps (Lu et al., 2024). Earlier studies on question answering demonstrate that an effective way to expand a query is to extract answer patterns or select terms that could be possible answers as expansion terms. Recently, some generation-augmented retrieval methods (Mao et al., 2021; Chen et al., 2022) focus on exploiting the knowledge captured in pre-trained language models (PLMs) (Roberts et al., 2020; Brown et al., 2020) to generate the potential answer as expansion terms.
Inspired by benchmark datasets containing detailed descriptions provided by human annotators, Zhang et al. (2020) propose the Query-to-Intent-Description (Q2ID) task that aims to generate a natural language intent description based on both relevant and irrelevant documents of a given query. Unlike their method, we directly model a query-aware irrelevant intent space in transformer models, leading to a more precise intent description.
2.2. Query-Focused Summarization
Query-focused summarization (QFS) is a subtask within text summarization that aims to generate a summary of a given text based on a specific query. Traditionally, the predominant approaches are unsupervised extraction methods (Wan and Xiao, 2009; Feigenblat et al., 2017) that generate summaries by ranking and extracting textual segments according to their similarity to other segments and their relevance to the query. In the past years, numerous QFS datasets (Kulkarni et al., 2020; Fabbri et al., 2022; Zhong et al., 2021) have been introduced, primarily based on answer selection or summaries for specific queries, leading to the development of many recent QA-motivated methods (Su et al., 2020, 2021; Egonmwan et al., 2019). Other works focus on query modeling through approaches such as the learning evidence ranking model (Xu and Lapata, 2021), jointly optimizing a latent query model and a conditional language model (Xu and Lapata, 2022), or designing pipeline models for abstractive summarizations. For instance, (Xu and Lapata, 2020) introduce a coarse-to-fine pipeline consisting of three estimators (relevance, evidence, and centrality estimators) that gradually discard less relevant segments.
Effectively extracting key insights from long documents has become a critical challenge. Vig et al. (2022) evaluate a two-step extract-then-abstract approach, where a scoring model is trained to select relevant portions of the text based on a query, which are then used as input to an abstraction model to generate a final summary. Additionally, they propose end-to-end methods that handle longer input texts by employing sparse attention mechanisms. Additional approaches focus on the use of a question-driven pretraining objective (Pagnoni et al., 2023), contrastive learning based on the scored segments (Sotudeh and Goharian, 2023), and joint modeling of token and utterance based on query-utterance attention (Liu et al., 2023a). In this work, we use state-of-the-art QFS methods as baselines for our proposed model.
3. Methods
Figure 2. Overview of our proposed pipeline. From left to right, we show (a) Intent-Driven Negative Augmentation method, (b) Contrastive decoder structure with dual cross-attention layers, (c) and (d) Contrastive learning via dual space modeling.
[…]
In this section, we present our approach QUIDS that generates descriptive query intents via dual space modeling. Figure 2 shows an overview. We first introduce the pipeline framework in section 3.1, then the data augmentation method to select intent-driven hard negative samples in section 3.2, and finally the contrastive mechanism via dual space modeling in section 3.3.
3.1. Pipeline Framework
We define the contrastive generation task as follows. Given a dataset $\mathcal{D}=\{(q,\mathcal{R},\mathcal{I},y)_{j}\}$ with $L$ samples, where $j\in\{0,1,...,L\}$: $q$ is a query, =$\mathcal{R}=\{r_{1},r_{2},...,r_{|\mathcal{R}|}\}$ is a collection of relevant documents for the query, =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ is a collection of irrelevant documents and $y$ is the ground truth query intent (i.e. the expected output). The goal is to learn the distinctions between relevant and irrelevant inputs based on a query while generating an intent description that exclusively highlights the relevant aspects related to the query. To achieve this, our training pipeline consists of 2 steps: (1) Intent-Driven Negative Augmentation (IDNA): mining hard negative documents as irrelevant documents from the entire dataset $D$ based on the query, its relevant document collections, and the ground truth intent, i.e.,
$IDNA(q,\mathcal{R},y,\mathcal{D})=\mathcal{I}^{\prime}$
where =$\mathcal{I}^{\prime}=\{i^{\prime}_{1},i^{\prime}_{2},...,i^{\prime}_{h}\}$ with $h$ the expected number of irrelevant documents. (2) Dual Space Modeling (DSM): contrastively generating a descriptive intent for the query by modeling query-aware relevant and irrelevant intent spaces, i.e.,
$DSM(q,\mathcal{R},i^{\prime})=\hat{y}$
3.2. Intent-Driven Negative Augmentation
Inspired by Liu et al. (2022) on choosing in-context sample strategies for in-context learning, we design a method to choose intent-aware hard negative samples based on semantic similarity. Specifically, we use a Sentence Transformer model (Reimers and Gurevych, 2019) to represent both positive and negative samples from the training data in a vector space. Then we choose negative samples close to the positive ones in this space. The negative sample augmentation makes the task more challenging for the model, hence improving its discriminative capabilities. As shown in Figure 2 (a), we augment hard negative samples using the following steps: For each query $q$ in the training set,
  1. (1)
    We rank its relevant documents $\mathcal{R}$ based on their cosine similarity to the concatenated embedding of the query and its ground truth intent $(q;y)$ where (;) denotes concatenation, resulting in $\mathcal{R^{*}}$.
  2. (2)
    We concatenate the ranked relevant documents $\mathcal{R^{*}}$ into a single document $R^{*}$ and obtain its embedding $h_{R^{*}}$.
  3. (3)
    The original irrelevant documents are initially ranked based on their cosine similarities to the newly concatenated relevant embedding $h_{R^{*}}$.
  4. (4)
    If the number of irrelevant documents is smaller than an expected size, the irrelevant document collection $\mathcal{I}^{\prime}$ is augmented by selecting hard negative samples from all other documents in the whole dataset $\mathcal{D}$ according to their cosine similarities to the new embedding $h_{R^{*}}$.
For step 1, 3 and 4, we use a pre-trained Sentence Transformers model (Reimers and Gurevych, 2019) trained on the MSMARCO Passage Ranking Dataset (Nguyen et al., 2016) to encode the documents. In practice, during step 4, we select documents whose cosine similarities exceed a predefined threshold 0.8 until the expected size is reached.
3.3. Dual Space Modeling
We use a Transformer-based encoder-decoder architecture, with the BART-large model (Lewis et al., 2020) and the T5-large model (Raffel et al., 2020a)as the backbone, as illustrated in Figure 2 (b). To capture the relationship between a query and its relevant and irrelevant documents, we adopt a Siamese dual encoder architecture. Based on the encoder outputs, contrastive learning is performed in the representation space to differentiate between embeddings for relevant and irrelevant documents. Correspondingly, we design a contrastive decoder to model query-aware relevant and irrelevant intent spaces in a disentangled manner.
3.3.1. Representation space modeling
Naturally, we expect the encoder to model and distinguish the relevant documents and irrelevant documents concerning a query in a feature representation space. To this end, we implement a dual encoder architecture, with each encoder being a cross-encoder, so that each cross-encoder directly models the relevance of a relevant or irrelevant document given a query by jointly encoding the query-document pair.
Intuitively, relevant documents are assumed to share similar topics, reflecting their connection to the specific information need behind a query to varying degrees. Therefore, it is natural to encode relevant documents collectively by concatenating them according to the ranking outlined in Section 3.2, Step 1. In contrast, irrelevant documents can be irrelevant to a query in diverse ways, making it impractical to model a meaningful and comprehensive irrelevant feature representation space. Instead, we focus on a single irrelevant document $i^{\prime}$ at each training step, using a hard negative sample from the augmented irrelevant document collection $\mathcal{I^{\prime}}$. We aim to model a feature representation space for relevant documents, reflecting their shared topics, while ensuring the irrelevant feature representation space is distant from the relevant one due to the dissimilarity of their topics. To achieve this, we project both of the encoder last hidden states into a feature representation space via a linear layer. The feature representation for each relevant document is obtained by applying average pooling over the token representations corresponding to the document. We then pull the relevant feature representations closer together while pushing the irrelevant representation further away from each relevant one (Figure 2 (c)). The optimization objective of the relevant feature representation space is to minimize:
(1) $\mathcal{L}_{rel}=\sum^{k}_{m=1}\sum^{k}_{n=m+1}d(e_{m},e_{n})$
where $e$ is the embedding of each relevant document, $k$ is the number of relevant documents, and $d$ is a distance function. We use cosine distance for $d$ in this work. For irrelevant feature representation space, we optimize the margin loss function:
(2) $\mathcal{L}_{irrel}=\sum^{k}_{m=1}max(t-d(e_{m},\bar{e}),0)$
where $\bar{e}$ is the embedding of the irrelevant document, and $t$ is a margin parameter, set to 1 in our case. We combine the relevant and irrelevant loss to obtain the encoder contrastive loss as follows:
(3) $\mathcal{L}_{encoder}=\mathcal{L}_{rel}+\mathcal{L}_{irrel}$
3.3.2. Disentangling space modeling
In the previous encoding phase, we obtain contextual feature representations for relevant and irrelevant documents given a query. In the decoder part, we want to make sure the decoder generates intent descriptions based on these information sources. Hence, we design a contrastive decoder by adding an additional cross attention so that it can attend to both relevant and irrelevant documents. To effectively disentangle the contrastive information learned by the model during the decoding process, we apply additional contrastive learning in disentangling space. As shown in Figure 2 (d), the model is designed to capture the contrast between relevant and irrelevant information spaces, allowing it to model both relevant and irrelevant intent spaces. This enables the generation of more precise and nuanced intent descriptions by focusing on the relevant intent space while minimizing attention on the irrelevant one.
Our decoder adopts a Transformer architecture, composed of N identical decoder layers. In the $l$-th decoder layer, at the $z$-th decoding step, we obtain hidden states $h_{self,z}^{l}$ by employing masked self-attention layers, to make sure the prediction of position $z$ depends only on the predictions before $z$. Based on $h_{self,z}^{l}$, we compute relevant document hidden states $h_{rel,z}^{l}$ by applying multi-head attention with cross-attention (MHAtt) to relevant encoder output:
(4) $h_{rel,z}^{l}=MHAtt(h_{self,z}^{l},h_{R^{*}})$
Similarly, we get the irrelevant document hidden states by attending to irrelevant encoder output:
(5) $h_{irrel,z}^{l}=MHAtt(h_{self,z}^{l},h_{i^{\prime}})$
From preliminary results, we found that a simple linear combination of $h_{self,z}^{l}$, $h_{rel,z}^{l}$, and $h_{irrel,z}^{l}$ works well to serve as the decoder hidden state to produce the distribution over the target vocabulary:
(6) $h_{combine,z}^{l}=h_{self,z}^{l}+h_{rel,z}^{l}-h_{irrel,z}^{l}$
(7) $P^{vocab}_{z}=Softmax(W(h_{combine,z}^{N}))$
where $W$ indicates a linear transformation. We optimize the model with the negative log likelihood (NLL) objective function to predict the target words:
(8) $\mathcal{L}_{NLL}=-\sum^{|y|}_{z=1}logP^{vocab}_{z}(y_{z})$
Corresponding to the representation space contrastive learning, we perform another contrastive learning in the newly proposed disentangling space using hidden states from the last decoder layer. We apply an additional linear layer to $h_{self}^{N}$, $h_{rel}^{N}$, and $h_{irrel}^{N}$, projecting them into a new representation space. We then obtain the embeddings $f_{c}$, $f_{r}$, $f_{i^{\prime}}$ by pooling these projected vector representations.
We follow the approach of SimCLR (Chen et al., 2020) and use from-batch negative samples $\mathcal{B}$ in the InfoNCE loss (He et al., 2020):
(9) $\mathcal{L}_{decoder}=-log\frac{exp(cos(f_{c},f_{r})/\tau)}{\sum_{i^{\prime}\in\mathcal{B}}exp(cos(f_{c},f_{i^{\prime}})/\tau)}$
where $\tau$ is the temperature and $cos(\textperiodcentered,\textperiodcentered)$ defines cosine similarity.
Finally, we combine the original NLL loss together with encoder and decoder loss to obtain the overall loss $\mathcal{L}$ to update all learnable parameters in an end-to-end learning setting:
(10) $\mathcal{L}_{NLL}=\lambda_{0}\mathcal{L}_{NLL}+\lambda_{1}\mathcal{L}_{Encoder}+\lambda_{2}\mathcal{L}_{Decoder}$
where $\lambda$ parameters control the balance between the three losses, with their total sum equal to 1.
4. Experimental Settings
4.1. Dataset
Q2ID (Zhang et al., 2020) is a benchmark dataset for query-to-intent description, derived from existing TREC and SemEval collections: TREC is an ongoing series of benchmark activities in IR. Q2ID comprises the TREC Dynamic Domain track 2015, 2016, 2017 on dynamic, exploratory search and the TREC 2004 Robust Track on consistency of retrieval technology. These tracks together contribute to 510 entries in the Q2ID benchmark dataset. SemEval is a series of workshops on Semantic Evaluation. Q2ID comprises the English SemEval-2015 Task3 and SemEval-2016 Task3 on Community Question Answering, contributing 4878 entries to the Q2ID dataset.
For each query in the collections, documents with multi-graded relevance labels were converted into binary labels, indicating whether they are relevant or not to the query. In total, the dataset comprises 5,358 entries, each formatted as a quadruple: ¡query, relevant documents, irrelevant documents, intent description¿, with intent descriptions being human-written narratives that provide detailed descriptions of the search intent behind the queries. The length statistics are an average of 7.2 tokens per query and 45.5 tokens per intent description. We split the 5338 queries into training (5000), validation (100), and test (258) query sets according to the original split setting of the Q2ID benchmark dataset.
4.2. Baselines
Since query-focused summarization task and query intent generation task are closely aligned in their focus on user queries and the extraction of relevant information, we evaluate our model against state-of-the-art models from the Q2ID benchmark, QFS task, and additionally include LLM zero-shot and two-shot baselines.
  • Pretrained sequence-to-sequence model baselines: T5 (Raffel et al., 2020b): a Transformer-based encoder-decoder (Vaswani et al., 2017) model trained on a diverse and extensive dataset. We use a pretrained T5-large model that we finetuned on the original Q2ID training dataset. BART (Lewis et al., 2020): also a transformer-based encoder-decoder model, trained by corrupting documents and then optimizing a reconstruction loss. The BART model serves as the backbone of our QUIDS model.
  • Query-to-intent description (Q2ID) baseline: CtrsGen (Zhang et al., 2020): a Q2ID model using a bi-directional GRU as encoder architecture. During decoding, it computes contrast scores by considering irrelevant documents to adjust sentence-level attention weights in the relevant documents.
  • Large Language Model (LLM) baseline: LLama3.1 (AI, 2024): We use the Llama3.1-8B instruction-tuned text-only model in both zero-shot and two-shot settings and conduct five experimental runs for each setting. For two-shot setting, we randomly using two different examples per run – one sourced from TREC and the other from SemEval.
  • Query Focused Summarization (QFS) baselines: RelReg (Vig et al., 2022) and RelRegTT (Vig et al., 2022): two-step approaches for QFS consisting of an score-and-rank extractor and an abstractor. The extractor is trained to predict ROUGE relevance scores and then the ranked results based on ROUGE are passed to the abstractor. SegEnc (Vig et al., 2022): an end-to-end approach tailored for handling longer input texts. SegEnc splits a long input into fixed-length overlapping segments and encodes them separately. The encoding sequences are concatenated so that the decoder can attend to all encoded segments jointly. Socratic (Pagnoni et al., 2023): an unsupervised, question-driven pretraining approach designed to tailor generic language models for controllable summarization tasks. Qontsum (Sotudeh and Goharian, 2023): an abstractive summarizer that applied Generative Information Retrieval (GIR) techniques. It builds on SegEnc by adding a segment scorer and contrastive learning modules.
We train the QFS baselines on the Q2ID dataset using the original code provided by the authors, except for Qontsum, which we independently reproduced.
4.3. Implementation Details
We build our method based on the BART-large (Lewis et al., 2020) model using Huggingface Transformers (Wolf et al., 2019). CrossRel attention and CrossIrrel attention layers in the decoder blocks share the same initial weights from the pre-trained BART-large model. For the IDNA step, we set the expected number of irrelevant documents per query to three in the training set. To ensure this, we augment 1,984 queries so that each query includes at least three irrelevant documents. We train the model for 10 epochs using the Adam optimizer with a learning rate of 0.0001. The final checkpoints are chosen based on the average ROUGE-{1, 2, L} scores against the ground truth query intent obtained from the validation set. During decoding, we set the maximum length to 256 tokens and apply beam search with a beam size of 4, using a no-repeat n-gram size of 3. We optimize $\lambda$ during training and fix on $(\lambda_{0}=0.2,\lambda_{1}=0.2,\lambda_{2}=0.6)$ to balance the three losses. All experiments are conducted five times, and we report the average results across these runs. The implementation details for the other baseline models can be found in Appendix A.1.
4.4. Evaluation Metrics
We conduct three kinds of evaluations using different evaluator resources: automatic evaluation, LLM-based evaluation and human evaluation. For automatic evaluation, we report recall scores on ROUGE-{1, 2, L} following Zhang et al. (2020), along with BERTScore (Zhang et al., 2019), which assesses semantic and syntactic similarity beyond exact word matches. Additionally, we conduct a human evaluation study using 50 randomly selected test samples. Three PhD students in Computer Science scored intent descriptions from our model and the best baseline, without knowing which model produced them. They rated both models on four customized qualitative criteria, with scores ranging from 1 (worst) to 5 (best), and the final scores were averaged across the three annotators.
  • Fluency: to what extent the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
  • Factual Alignment: to what extent the generated query intent description is factually aligned with the ground truth intent.
  • Inclusion score: how well the generated query intent includes important details from the query and relevant documents.
  • Exclusion score: how well the generated query intent description excludes information present in the irrelevant documents that is not relevant to the query and relevant documents.
However, conventional automatic evaluation metrics tend to show weak correlation with human judgments. Recent study (Fu et al., 2024; Liu et al., 2023b) propose using LLMs as evaluators for natural language generation (NLG) systems, scoring generated outputs based on generation probability rather than reference targets. Following the method of (Liu et al., 2023b), we use LLaMa3.1-8B [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ as instruction-tuned evaluators to assess the generated intent across four qualitative metrics. Specifically, we define the evaluation task and criteria, prompting the LLM to generate chain-of-thoughts (CoT) for the ‘Evaluation Steps’. For LLaMa3.1-8B, we use the output token probabilities from the LLMs to normalize the scores and take their weighted summation as the final results: $score=\sum_{i=1}^{n}p(s_{i})\times s_{i}$ where $S=\{s_{1},s_{2},...,s_{n}\}$ represents the predefined score set from the prompt, with a maximum value of 5 in our case. For the close-sourced GPT-4o, we sample 20 times to estimate the token probabilities. Example prompts are in Appendix B.
5. Experimental Results
Table 1. Performance between our model and baselines in terms of automatic evaluation (%). $\dagger$ indicates reported performance from previous work. ‘-’ means the result is inaccessible. * indicates the model outperforms the best baseline significantly with paired t-test at $p$-value $<$ 0.05 level. The best results are highlighted in bold, while the best baseline results are underlined.
Models RG-1 RG-2 RG-L BS
[…]
T5-large 28.87 13.91 23.85 61.64
BART-large 30.70 13.91 24.63 62.07
LLaMa3.1 zero-shot 29.28 7.42 20.90 57.26
LLaMa3.1 2-shot 32.75 9.54 24.34 57.89
RelReg 26.67 12.83 21.99 59.24
RelRegTT 27.21 12.77 22.25 59.60
SegEnc 31.83 14.29 25.18 62.15
+ SOCRATIC Pret. 31.38 13.88 24.91 62.26
QONTSUM 31.18 14.26 24.87 62.03
QUIDS_T5 29.40 13.95 24.23 62.00
QUIDS_BART 34.47* 14.86* 26.77* 63.55*
5.1. Overall Results
We compare model performance between QUIDS and baselines in Table  1. The results show that: (1) Our model QUIDS outperforms all other baselines; (2) Our approach is compatible with both T5 and BART architectures, which both have a transformer-based encoder-decoder structure. This demonstrates the effectiveness of our design in the query intent generation task. Notably, BART-large outperforms T5-large despite having nearly half the model size, highlighting its efficiency in this task. (3) We implemented QFS models for the query intent generation task, which already outperform the baseline model, CtrsGen. However, our design further improves ROUGE-1 by 2.64 and ROUGE-L by 1.59 compared to the best QFS model, SegEnc; (4) In our experiments, the two-shot setting with the LLaMa3.1 8B model significantly outperforms the zero-shot setting in ROUGE scores, while showing only minor improvements in BERTScore. This suggests that without fine-tuning, intent generation may be lexically similar but semantically misaligned.
5.2. Ablation Study
Table 2. Ablation study of our QUIDS model with its variants under automatic evaluation (%).
Models RG-1 RG-2 RG-L BS
QUIDS w/o IDNA 33.48 14.20 25.95 63.17
QUIDS w/o RSM 34.57 14.39 26.38 63.62
QUIDS w/o DSM 33.45 13.46 25.88 63.33
QUIDS 35.95 14.80 27.21 64.33
We conduct an ablation study on the test set to assess the impact of various modules in our proposed approach under three conditions: (1) We remove the IDNA module to examine the effect of data preprocessing (w/o IDNA); (2) We remove contrastive learning from the encoder and (3) decoder separately to assess the importance of representation space modeling (w/o RSM) and disentangling space modeling (w/o DSM), respectively. Table 2 shows the results. Removing IDNA leads to a 1.16-point drop in BERTScore, suggesting that incorporating harder irrelevant documents helps the model capture relevant intent information more effectively. Removing contrastive learning from the decoder leads to a significant drop in all metrics, highlighting its importance in modeling a distinguishable intent space and generating accurate intent descriptions. Finally, removing contrastive learning from the encoder results in a slight decrease, indicating that representation space modeling aids the model in distinguishing relevant from irrelevant information. Due to the similar architecture and experimental performance, we omit the T5 model results.
5.3. Human and LLM Evaluation
We analyze the quality of the generated intents by our model QUIDS and the best baseline model SegEnc with human evaluation and LLM-based approaches. Human inter-annotator agreement is assessed with weighted Cohen’s kappa, revealing fair to moderate consistency across the four metrics. As shown in table 3, QUIDS consistently receives higher scores from all three evaluators, except for Factual Alignment, where human evaluators prefer SegEnc’s outputs.
Table 3. Comparison of human evaluation and LLM evalaution in terms of Fluency (Fluen.), Factual Alignment (Align.), Inclusion score (Inclu.) and Exclusion score (Exclus.).
Method Model Fluen. Align. Inclu. Exluc.
Human SegEnc 4.71 3.70 3.88 4.71
QUIDS 4.73 3.61 3.97 4.73
GPT-4o SegEnc 3.70 2.64 2.67 3.83
QUIDS 3.71 2.79 2.69 4.00
LLaMa3.1 SegEnc 3.25 2.91 3.17 4.07
QUIDS 3.48 3.17 3.42 4.11
Figure 3. Distribution of human and LLM evaluation sores on four qualitative metrics.
[…]
Figure 4. (a) Visualizes the token-level decoder cross-attention weights for a relevant document snippet when generating the token ‘impacted’ in Intent 1, without an irrelevant document provided. (b) for relevant document and irrelevant document when generating the token ‘effects’ in Intent 2, with an irrelevant document provided. Deeper color indicates a higher value.
[…]
To examine the performance differences between humans and LLMs in this task, we present boxplots in Figure 3 that display the score distributions for each metric. Additionally, we calculate the Spearman and Kendall-Tau correlations between the evaluations from LLMs and human judgments, as shown in Table 5 in Appendix A.2. From this analysis, we derive the following insights:
  • Humans tend to assign higher scores than GPT-4o and LLaMa 3.1 across all metrics, particularly for Fluency and Exclusion, where scores nearly reach 5. The variability in fluency scores suggests humans are more tolerant of less fluent text.
  • Human evaluations exhibit the narrowest range, while GPT-4o shows a similar range but with more critical scores. As supported by the correlation table in Appendix A.2, GPT-4o aligns more closely with human evaluations, particularly in Factual Alignment. However, a difference in model preferences for this metric prompts further analysis of sub-dataset differences in Section 5.4.
  • For the metrics Factual Alignment and Inclusion, all methods tend to exhibit a wider distribution with generally lower scores. This suggests variability in the assessments of these two metrics, potentially indicating inconsistency in how both human and LLM evaluators interpret the scoring criteria.
5.4. Analysis of Intent Types
For analysis of the effect of intent type, we classify the queries according to their underlying search intent into two categories: (1) Informational Intent: Natural language questions seeking detailed information or solutions, typically longer and contextual. Queries from the SemEval dataset fall under this category. (2) Exploratory Intent: Term-based queries aimed at broad exploration with minimal context or structure. Queries from the TREC datasets are categorized here. In the 258 test samples, there are 20 queries with exploratory intents and 238 with informational intents. As shown in Table 4, queries with exploratory intents substantially outperform those with informational intents, achieving 60% higher ROUGE-2 scores and 20% higher BERTScores. This indicates that our model is better suited for exploratory queries. This finding contrasts with the results of (Zhang et al., 2020), where the CtrsGen model performed slightly better on the informational SemEval queries than on the exploratory TREC queries. A potential explanation is that the backbone language model used in our approach more freely generates text than the GRU model used in (Zhang et al., 2020), particularly when reconstructing complex scenarios for informational intents. This is an aspect that makes our approach more suitable for exploratory search rather than informative search. We additionally conduct a breakdown analysis by intent types for the human and LLM results in Appendix A.3.
Table 4. Comparison of automatic evaluation on our model QUIDS for different intent types.
Intent RG-1 RG-2 RG-L BS
Informational 35.69 14.51 26.82 63.88
Exploratory 41.55 23.24 38.28 76.65
5.5. Analysis of Document Length
In real-world scenarios, relevant documents associated with a query can vary in length. To assess the impact of document length on the generated output, we categorize the documents into three groups based on the BART-large input limit (1024 tokens): documents shorter than 512 tokens are classified as short documents; those ranging from 512 to 1024 tokens are termed medium-sized documents; and documents longer than 1024 tokens, which the model will truncate, are labeled as long documents. In our test dataset, there are 204 short, 25 medium, and 29 long documents. The overall differences in performance among all the input length categories are minimal across all metrics (also see Figure 6 in Appendix A.4). This indicates that 1024 tokens provide sufficient information for our model to effectively capture the underlying query intent. This may explain why QFS approaches, particularly those designed to handle lengthy documents, do not demonstrate significant advantages in the query intent generation task.
5.6. Case Study
To intuitively understand how irrelevant documents influence our model, we compare the generated intents for a query ‘Freon-12’ from test set with and without providing an irrelevant document. Figure 4 illustrates the token-level decoder cross-attention weights for both a relevant document snippet, and an irrelevant document snippet when provided. As shown in Figure 4 (a), without an irrelevant document provided, when generating the word ‘impacted’, the model’s focus is centered on the economic effects on cars, as indicated by the mention of ‘price’ in Intent 1. With an irrelevant document provided in Figure 4 (b), when generating the word ‘effects’ in Intent 2, similar attentions can be detected on ‘prices’, ‘\$’, ‘costs’, ‘auto’ for both relevant and irrelevant documents. However, as shown in Generated Intent 2, our model now treats tokens related to prices and cars in relevant documents as irrelevant. This demonstrates that out model effectively generates intent descriptions that excludes this unrelated information.
6. Conclusion
In this paper, we introduced a novel dual-space modeling approach for the task of query intent generation. Our approach implements contrastive learning in both the encoding and decoding phases, combined with intent-driven hard negative augmentation during data preprocessing, to automatically generate detailed and precise intent descriptions. Experimental results show that our model effectively filters out irrelevant information from the relevant intent space, leading to more accurate intent descriptions and greater transparency in the retrieval process. For future work, we aim to extend our approach to conversational search by first mining exploratory needs and then explaining the understanding of query intents.
Acknowledgements.
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
Appendix A Appendix A
A.1. Implementation details for baseline models
RelReg and RelRegTT share the same abstractor, a BART-large model, which also serves as the backbone model for SegEnc, Socratic, and Qontsum. For RelReg and RelRegTT, we use an input segment length of 1024, whereas SegEnc-based models utilize an input segment length of 512, with a total input length of 4096. For Socratic training, we use the checkpoint pretrained on Books3 (Csaky and Recski, 2020) from the Huggingface Model Hub [^3] https://huggingface.co/Salesforce/socratic-books-30M and fine-tune it on Q2ID dataset using SegEnc mechanism. We reproduce the work of Qontsum with the segment length of 512 tokens, temperature of 0.6 and ($\lambda_{0}=0.6,\lambda_{1}=0.2,\lambda_{2}=0.2$) in joint learning. For all models that divide input text into segments, we apply a 50% overlap between each segment and its adjacent one.
A.2. Correlation with Human Evaluation
Table 5. Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations between Human evaluation and LLM evaluation of different metrics.
Correlation Model Fluen. Align. Inclu. Exluc.
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Corr (Human, GPT-4o) SegEnc 0.377 0.304 0.409 0.310 0.184 0.159 0.369 0.306
QUIDS 0.476 0.375 0.517 0.412 0.252 0.199 0.361 0.286
Corr (Human, LLaMa3.1) SegEnc 0.243 0.187 0.459 0.344 0.143 0.111 0.224 0.177
QUIDS 0.343 0.278 0.553 0.423 0.325 0.248 0.158 0.126
We assess the correlation between human and LLM evaluators across four qualitative evaluation criteria, presenting the Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations for the best SOTA model SegEnc and our QUIDS model in Table 5. Overall, our model demonstrates significantly higher human correspondence across all metrics compared to SegEnc, with the exception of the Exclusion score. Correlation performance varies by metric; for Fluency and Factual Alignment—criteria requiring less contextual information—there is a relatively higher degree of agreement with human evaluations. In contrast, the Inclusion and Exclusion scores, which depend on diverse and contextual sources, show lower correlation, suggesting that humans and LLM evaluators adopt different evaluation strategies for more complex criteria. Additionally, we observe that different LLM evaluators exhibit human-like evaluation behaviors across various metrics: LLaMa3.1 shows greater human correspondence in Factual Alignment and Inclusion scores, whereas GPT-4o aligns more closely with human evaluations in Fluency and Exclusion scores.
A.3. Breakdown Analysis on Intent Types
Table 6. Comparision of Human and LLM evaluation on informational and exploratory intent types on our model.
Method Intent Fluen. Align. Inclu. Exluc.
Human Info. 4.71 3.51 3.98 4.73
Expl. 4.85 4.04 3.88 4.74
GPT-4o Info. 3.64 2.70 2.63 3.95
Expl. 4.02 3.17 2.96 4.20
LLaMa3.1 Info. 3.33 2.96 3.35 3.99
Expl. 3.72 4.11 3.72 4.64
Figure 5. Boxplot of evaluation scores on 4 metrics of our model on different intent types.
[…]
In Table 3, we observe a human preference over the SegEnc model on metric Factual Alignment, which measures how well the generated query intent description is factually aligned with the ground truth intent. We guess it is due to the model performance difference on different sub-datasets, or on different intent types. And hence we further analyse the human evaluations on different intent types. Table 6 present human and LLM evaluations on our model regarding two intent types. While humans prefer our model for exploratory intent with 4.04 (QUIDS) vs. 3.81 (SegEnc), SegEnc is favored for informative intent with 3.51 (QUIDS) vs. 3.68 (SegEnc). Since informative intent queries dominate, this leads to a lower average score for our model. We also present the evaluation scores distribution on different intent types over 50 the same test samples in subsection 5.4. Figure 5 is a breakdown of Figure 3 on our model for four metrics. Based on Table 6 and Figure 5, we have following insights:
  • Exploratory intent types generally outperform informational intent types across all metrics, except for human evaluations on inclusion. This finding, based on 50 test samples from both human and LLM evaluations, is consistent with the automatic evaluation results for the full test dataset, as shown in Table 4.
  • The boxplots for informational intent on the right side of each metric in Figure 5 exhibit a distribution similar to the overall performance of our model presented in Figure 3, suggesting that informational intent queries dominate the dataset and largely drive performance outcomes. However, exploratory intent queries outperform in this particular task, indicating their stronger performance despite being less frequent.
A.4. Breakdown Analysis by document length
Figure 6 shows the results of the automatic evaluation, broken down based on document length.
Figure 6. Comparison of automatic evaluation on our model QUIDS on different input lengths.
[…]
Appendix B Appendix B
B.1. Evaluation prompt on Fluency
You will be given a query, relevant and irrelevant documents with respect to the query. You will also be given a generated query intent description based on the query and documents. The ground truth query intent description will also be provided.
Your task is to rate the query intent description on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
Evaluation Criteria:
Fluency Score (1-5) - This metric measures if the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
Evaluation steps:
1. Carefully review the provided query, relevant, and irrelevant documents to understand the context and content.
2. Read the ground truth query intent description to understand the ideal response. This serves as a benchmark for evaluating the fluency of the generated description.
3. Carefully read the generated query intent description. Focus on the fluency aspect, considering factors such as grammatical correctness, naturalness, clarity, coherence, and readability.
Assign a rating from 1 to 5 based on the level of fluency.
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}
Evaluation Form (scores ONLY):
- Fluency:
B.2. Evaluation prompt on Factual Alignment
You will be given a query, relevant and irrelevant documents with respect to the query. You will also be given a generated query intent description based on the query and documents. The ground truth query intent description will also be provided.
Your task is to rate the query intent description on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
Evaluation Criteria:
Factual Alignment (1-5) - This metric measures if the generated query intent description is factually aligned with the ground truth intent. Ensuring the facts presented in the generated description are correct and match those in the ground truth description. Verifying that all key facts and points mentioned in the ground truth are covered in the generated description without omission. Any hallucination that diverges from the ground truth should be flagged.
Evaluation steps:
1. Review the ground truth intent description for the central facts and points that convey the query’s purpose.
2. Read the generated intent description and list the main facts and points it conveys.
3. Compare the lists from the ground truth and generated descriptions for consistency in content. Look for alignment in terms of content, completeness, and accuracy.
4. Identify any key facts or points from the ground truth that are missing in the generated description (omissions) and note any information in the generated description that is not present or diverges from the ground truth (hallucinations).
Assign a rating from 1 to 5 based on the level of factual alignment.
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}
Evaluation Form (scores ONLY):
- Factual Alignment:
B.3. Evaluation prompt on Inclusion Score
You will be given a query, relevant and irrelevant documents with respect to the query. You will also be given a generated query intent description based on the query and documents. The ground truth query intent description will also be provided.
Your task is to rate the query intent description on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
Evaluation Criteria:
Inclusion Score (1-5) - This metric measures how well the generated query intent includes important details from the query and relevant documents. Assessing whether the generated description captures key elements that are directly relevant to the query. Evaluating if the generated description thoroughly includes significant points from the relevant documents. Ensuring that the included details are integrated in a way that maintains the context and importance as presented in the relevant documents.
Evaluation steps:
1. Review the query and relevant documents to extract the main facts, significant points, and key elements that directly address the query.
2. Read the generated query intent description and list the key details it includes.
3. Compare the key details and elements from the generated description with those identified from the query and relevant documents, checking for inclusion and alignment.
4. Assess how well the included details are integrated into the generated description, ensuring they maintain the context and importance as presented in the relevant documents.
Assign a rating from 1 to 5 based on the thoroughness and relevance of the included details.
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}
Evaluation Form (scores ONLY):
- Inclusion Score:
B.4. Evaluation prompt on Exclusion Score
You will be given a query, relevant and irrelevant documents with respect to the query. You will also be given a generated query intent description based on the query and documents. The ground truth query intent description will also be provided.
Your task is to rate the query intent description on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
Evaluation Criteria:
Exclusion Score (1-5) - This metric measures if the generated query intent description excludes information present in the irrelevant documents that is not relevant to the query and relevant documents. Evaluating whether the description effectively filters out information that is irrelevant to the query. Ensuring that the description does not include misleading or incorrect information found in the irrelevant documents. Evaluating whether the description effectively filters out information present in the irrelevant documents but focus on topics different from those in relevant documents.
Evaluation steps:
1. Carefully read through the irrelevant documents to pinpoint details, facts, or topics that are not relevant to the query and relevant documents.
2. Read the generated query intent description and extract the key details and points included in the description.
3. Compare the extracted content from the generated description with the irrelevant information identified in the irrelevant documents to check for the presence of any irrelevant details.
4. Assess how effectively the generated description filters out irrelevant information, ensuring it focuses only on the query and relevant documents.
Assign a rating from 1 to 5 based on the level of exclusion of irrelevant details.
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}
Evaluation Form (scores ONLY):
- Exclusion Score:

### v3

@@
-QUIDS: Query Intent Generation via Dual Space Modeling
-Abstract.
-Query understanding is a crucial component of Information Retrieval (IR), aimed at identifying the underlying search intent of textual queries.
-However, most existing approaches oversimplify this task into query classification or clustering, which fails to fully capture the nuanced intent behind the query.
-In this paper, we address the task of query intent generation: to automatically generate detailed and precise intent descriptions for search queries using relevant and irrelevant documents given a query.
-These intent descriptions can help users understand why the search engine considered the top-ranked documents relevant, and provide more transparency to the retrieval process.
-We propose a dual-space model that uses semantic relevance and irrelevance information in the returned documents to explain the understanding of the query intent.
-Specifically, in the encoding process, we project, separate, and distinguish relevant and irrelevant documents in the representation space.
-Then, we introduce a semantic decoupling model in the novel disentangling space, where the semantics of irrelevant information are removed from the relevant space, ensuring that only the essential and relevant intent is captured.
-This process refines the understanding of the query and provides more accurate explanations for the search results.
-Experiments on benchmark data demonstrate that our methods produce high-quality query intent descriptions, outperforming existing methods for this task, as well as state-of-the-art query-based summarization methods.
-A token-level visualization of attention scores reveals that our model effectively reduces the focus on irrelevant intent topics.
-Our findings open up promising research and application directions for query intent generation, particularly in exploratory search.
-Information Retrieval, Query intent generation, Contrastive learning, Transformers
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Users and interactive retrievalccs: Information systems Query intentccs: Computing methodologies Natural language generation
-1. Introduction
-Query understanding is a focal point in the Information Retrieval (IR) community, aimed at determining the underlying search intent of a textual query input.
-This facet of IR is crucial for several reasons.
-Firstly, it enables search engines to distinguish and filter out unwanted interpretations of the current search query, ensuring relevance and accuracy in results.
-This is important because queries tend to be short and underspecified and are therefore often ambiguous (Guo and Lan, 2020).
-Secondly, users benefit when search engines provide feedback, such as query expansions or more descriptive query intents (Chen et al., 2021).
-This enhances the search experience by enabling users to refine and adjust their queries based on insights derived from the provided query understanding.
-Figure 1.
-An application of the query intent generation task as a form of explanation to the user in exploratory search.
-The top-ranked documents are used as pseudo-relevant documents to generate the intent.
-Information relevant to forming the intent is highlighted in boldface.
-Existing methods for query understanding can be broadly categorized into three groups.
-Query classification methods (Broder, 2002; Verberne et al., 2013) aim to categorize queries into distinct classes based on intent taxonomies.
-However, many of these methods tend to abstract intents into broader categories, lacking granularity in their analysis.
-Second, query clustering methods (Wen et al., 2002; Hong et al., 2016) aim to group query intents into clusters, yet encounter challenges in precisely interpreting individual queries within these clusters.
-The third approach, which is the most effective way to capture query intent, is query intent description generation, where the intent is described in textual format.
-However, this task has been largely overlooked in existing work, with the most recent contribution by (Zhang et al., 2020), which still relies on RNN-based methods and does not leverage the advancements in transformer models.
-Figure 1 illustrates an example of query intent generation for the query ‘Freon-12’, which has an exploratory intent – aimed at exploring topics without a clearly defined goal  (Palagi et al., 2017).
-Since the true intent behind the query is unknown to the search engine, it relies on the retrieved documents to infer it.
-Thus, it treats the top-ranked documents as pseudo-relevant and the low-ranked documents as irrelevant.
-The intent description generated from both sets of documents is highlighted with a green background, while the information directly contributing to intent formation is marked in boldface.
-From the relevant documents, key terms like ‘Freon-12’, ‘refrigerant’, and ‘phased out’ are captured and emphasized in the intent description.
-In contrast, topics such as ‘costs’ and ‘prices’, which appear in both relevant and irrelevant documents, are deprioritized and excluded from the final intent description.
-The resulting intent generation illustrates how the search engine interprets the user’s query based on the retrieved documents.
-Furthermore, this approach allows users to reflect on their interests more clearly, potentially guiding them to reformulate their query more precisely and achieve more targeted search results.
-This iterative process of query refinement helps bridge the gap between exploratory search and more specific information needs, enhancing the overall search experience.
-In scenarios where the ground truth intent is available, a related task to train models for generating query intent is query-focused summarization (QFS).
-This task is to generate summaries tailored to specific user queries or information needs, drawing from pseudo-relevant (high-ranked) documents.
-However, the focus in QFS is on summarizing content rather than explicitly identifying or reproducing query intents.
-Typically, QFS does not involve irrelevant information, which means the resulting summaries may include unfiltered or irrelevant details that are not truly reflective of the query’s intent.
-On the other hand, previous work (Zhang et al., 2020) on the Q2ID task incorporates irrelevant documents for contrast, but it does not establish a direct connection between the query and irrelevant documents.
-Our approach explicitly models both relevant and irrelevant intent spaces in a query-aware manner.
-Specifically, we construct the final intent space by subtracting the irrelevant intent space from the relevant one.
-This subtraction ensures that the final intent description is closely aligned with the query while effectively filtering out irrelevant information.
-We propose a novel pipeline, implementing contrastive learning through dual space modeling in transformer-based (Vaswani et al., 2017) models, particularly T5 (Raffel et al., 2020a) and BART (Lewis et al., 2020).
-Specifically, we use a dual-encoder architecture and perform contrastive learning in the representation space using encoder embeddings of both relevant and irrelevant documents.
-During the decoding phase, we adapt the transformer decoder to conduct contrastive learning using decoder hidden states in the newly proposed disentangling space, where we model query-aware relevant and irrelevant intent spaces.
-This dual-space contrastive learning approach ensures that the model learns to differentiate between relevant and irrelevant documents, while also guiding it to produce high-quality intent descriptions with attention to the learned relevant and irrelevant intent space.
-Moreover, we introduce a method to augment hard negative documents driven by intent, to expand the irrelevant information representation space during training.
-We evaluate our model using the benchmark dataset Q2ID, sourced from TREC and SemEval, which includes two types of query intent.
-Our results demonstrate that our model outperforms state-of-the-art baselines in terms of ROUGE metrics and BERTScore.
-Additionally, we perform human evaluations and LLM-based evaluations using tailored metrics, and provide a comprehensive analysis of our approach.
-To further illustrate the effect of irrelevant documents, we conduct a case study in which we highlight the information used to generate a semantically rich intent token based on the decoder’s cross-attention weights at the token level for both relevant and irrelevant documents.
-In summary, our contributions are as follows:
-  • We introduce contrastive learning in both the representation space and the disentangling space of transformer models, effectively capturing contrasting information from relevant and irrelevant documents.
-  • Our model generates high-quality intent descriptions, with performance significantly enhanced by incorporating hard negative data augmentation during training.
-  • We perform a thorough evaluation of our model, providing us with insights into the model’s strengths and weaknesses, as well as its potential application scenarios.
-2. Related Work
-2.1. Query Understanding
-The most closely related tasks in query understanding include query classification, query clustering, and query expansion.
-Query classification (Guo and Lan, 2020; Li et al., 2023) aims to classify queries into one or more pre-defined target taxonomies, based on factors like intent (Broder, 2002; Rose and Levinson, 2004), topic (Li et al., 2005; Beitzel et al., 2005), information type (Pasca and Harabagiu, 2001) or other classification dimensions.
-However, due to the typically short and ambiguous nature of search queries, as well as their evolving meanings, these taxonomies may lack sufficient granularity and become outdated over time (Guo and Lan, 2020).
-This can lead to the loss of detailed information in the classification process.
-Query clustering attempts to discover search topics or intents by mining clusters of queries.
-The similarity between queries is typically measured using data from user click-through logs (Wen et al., 2002; Rani and Babu, 2019; Lu et al., 2024) or top-ranked documents (Hong et al., 2016).
-In a recent application, query clustering has been employed in search engines to display real-time trending user queries, capturing shifts in queries and news articles over consecutive timestamps (Lu et al., 2024).
-Earlier studies on question answering demonstrate that an effective way to expand a query is to extract answer patterns or select terms that could be possible answers as expansion terms.
-Recently, some generation-augmented retrieval methods (Mao et al., 2021; Chen et al., 2022) focus on exploiting the knowledge captured in pre-trained language models (PLMs) (Roberts et al., 2020; Brown et al., 2020) to generate the potential answer as expansion terms.
-Inspired by benchmark datasets containing detailed descriptions provided by human annotators, Zhang et al.
-(2020) propose the Query-to-Intent-Description (Q2ID) task that aims to generate a natural language intent description based on both relevant and irrelevant documents of a given query.
-Unlike their method, we directly model a query-aware irrelevant intent space in transformer models, leading to a more precise intent description.
-2.2. Query-Focused Summarization
-Query-focused summarization (QFS) is a subtask within text summarization that aims to generate a summary of a given text based on a specific query.
-Traditionally, the predominant approaches are unsupervised extraction methods (Wan and Xiao, 2009; Feigenblat et al., 2017) that generate summaries by ranking and extracting textual segments according to their similarity to other segments and their relevance to the query.
-In the past years, numerous QFS datasets (Kulkarni et al., 2020; Fabbri et al., 2022; Zhong et al., 2021) have been introduced, primarily based on answer selection or summaries for specific queries, leading to the development of many recent QA-motivated methods (Su et al., 2020, 2021; Egonmwan et al., 2019).
-Other works focus on query modeling through approaches such as the learning evidence ranking model (Xu and Lapata, 2021), jointly optimizing a latent query model and a conditional language model (Xu and Lapata, 2022), or designing pipeline models for abstractive summarizations.
-For instance, (Xu and Lapata, 2020) introduce a coarse-to-fine pipeline consisting of three estimators (relevance, evidence, and centrality estimators) that gradually discard less relevant segments.
-Effectively extracting key insights from long documents has become a critical challenge.
-Vig et al.
-(2022) evaluate a two-step extract-then-abstract approach, where a scoring model is trained to select relevant portions of the text based on a query, which are then used as input to an abstraction model to generate a final summary.
-Additionally, they propose end-to-end methods that handle longer input texts by employing sparse attention mechanisms.
-Additional approaches focus on the use of a question-driven pretraining objective (Pagnoni et al., 2023), contrastive learning based on the scored segments (Sotudeh and Goharian, 2023), and joint modeling of token and utterance based on query-utterance attention (Liu et al., 2023a).
-In this work, we use state-of-the-art QFS methods as baselines for our proposed model.
-3. Methods
-Figure 2.
-Overview of our proposed pipeline.
+QUIDS: Query Intent Description for Exploratory Search via Dual Space Modeling
+Abstract
+In exploratory search, users often submit vague queries to investigate unfamiliar topics, but receive limited feedback about how the search engine understood their input.
+This leads to a self-reinforcing cycle of mismatched results and trial-and-error reformulation.
+To address this, we study the task of generating user-facing natural language query intent descriptions that surface what the system likely inferred the query to mean, based on post-retrieval evidence.
+We propose QUIDS, a method that leverages dual-space contrastive learning to isolate intent-relevant information while suppressing irrelevant content.
+QUIDS combines a dual-encoder representation space with a disentangling decoder that works together to produce concise and accurate intent descriptions.
+Enhanced by intent-driven hard negative sampling, the model significantly outperforms state-of-the-art baselines across ROUGE, BERTScore, and human/LLM evaluations.
+Our qualitative analysis confirms QUIDS’ effectiveness in generating accurate intent descriptions for exploratory search.
+Our work contributes to improving the interaction between users and search engines by providing feedback to the user in exploratory search settings.
+[^1] Our code is available at https://github.com/menauwy/QUIDS
+QUIDS: Query Intent Description for Exploratory Search
+via Dual Space Modeling
+Yumeng Wang Leiden University Leiden, The Netherlands y.wang@liacs.leidenuniv.nl          Xiuying Chen MBZUAI Abu Dhabi, United Arab Emirates xiuying.chen@mbzuai.ac.ae          Suzan Verberne Leiden University Leiden, The Netherlands s.verberne@liacs.leidenuniv.nl
+1 Introduction
+Figure 1: A user-facing application of query intent generation in exploratory search.
+The system’s inferred intent is generated by contrasting top-ranked (pseudo-relevant) and low-ranked (pseudo-irrelevant) documents.
+Key information contributing to the inferred intent is shown in bold.
+In exploratory search (Palagi et al., 2017), users often issue vague or underspecified queries to investigate unfamiliar topics through iterative refinement.
+This process gives rise to a persistent usability challenge, which we call the dual-blind problem: Users are uncertain about how to express their information needs; as a result they formulate ambiguous queries; the system silently infers the user’s intent based on these ambiguous queries, without providing explicit feedback and retrieves mixed-quality results at best.
+This leads to a self-reinforcing cycle of the user receiving results that are misaligned with their actual information need.
+This cycle is difficult to break with traditional query understanding methods Li et al.
+(2023); Lu et al.
+(2024); Wang et al.
+(2023) that operate in the pre-retrieval stage.
+Their goal is to optimize ranking effectiveness, not to provide feedback to the user.
+They offer little transparency about how the system arrived at its results, which is especially a problem when users are unsure of their intent.
+In this work, we study the task of generating a natural-language query intent description (Zhang et al., 2020) that reflects what the system likely inferred the query to mean.
+The description is generated in the post-retrieval stage, incorporating system-inferred relevance of documents.
+These descriptions are not mechanistic explanations of the ranker, but instead serve as user-facing proxies of the system’s inferred intent.
+By contrasting top-ranked (pseudo-relevant) and low-ranked documents, the intent descriptions provide feedback that helps users identify mismatches between their intended and inferred query meanings.
+This feedback supports more effective query refinement and improves the overall search experience.
+Figure 1 illustrates an application scenario of query intent generation in exploratory search.
+Since the ground truth intent behind the query is unknown to the search system, it relies on the retrieved documents to infer the query intent.
+From the top-ranked documents (considered relevant by the search engine), key terms like ‘Freon-12’, ‘refrigerant’, and ‘phased out’ are captured and emphasized in the intent description.
+In contrast, topics such as ‘costs’ and ‘prices’, which appear in both high-ranked and low-ranked documents, are excluded from the final intent description.
+The resulting description provides diagnostic feedback on how the system understood the query, helping users assess whether the retrieved results align with their latent intent.
+We introduced a novel dual-space modeling approach, QUIDS, for the query intent generation task.
+It models query intent through dual-space contrastive learning by performing contrastive learning in two complementary spaces, explicitly separating intent-relevant and irrelevant semantic information.
+Specifically, the method consists of: (i) a representation space via dual encoders and (ii) a novel disentangling space in the decoder.
+This dual-space design enables the model to subtract irrelevant semantics from relevant ones, generating concise and more accurate intent descriptions.
+Furthermore, we propose an intent-driven hard negative sampling strategy to expand the irrelevant representation space and improve contrastive learning during training.
+Experiments on the Q2ID benchmark (Zhang et al., 2020), including TREC and SemEval datasets, show that our model significantly outperforms strong baselines, including the prior Q2ID-specific method, LLM-based, and Query-focused Summarization methods, both in automatic and human evaluations.
+Qualitative analysis confirms that QUIDS effectively filters out distracting or misleading content and generates concise intent descriptions.
+Our contributions are: (i) Our model generates high-quality intent descriptions, with performance significantly enhanced by incorporating hard negative data augmentation during training.
+(ii) We introduce contrastive learning in both the representation space and the disentangling space of transformer models, effectively capturing contrasting information from relevant and irrelevant documents.
+(iii) We perform a thorough evaluation of our model, providing us with insights into the model’s strengths and weaknesses, as well as its potential application scenarios, especially for exploratory search.
+2 Related Work
+2.1 Query Understanding
+Our work is related to traditional query understanding tasks such as classification Broder (2002); Verberne et al.
+(2013), clustering Wen et al.
+(2002); Hong et al.
+(2016), and expansion Wang et al.
+(2023); Mo et al.
+(2023); Jagerman et al.
+(2023).
+However, unlike these methods, which operate in the pre-retrieval stage to optimize retrieval effectiveness, Zhang et al.
+(2020) proposes the Query-to-Intent-Description (Q2ID) task in the post-retrieval stage that aims to generate search systems’ inferred intent of a user query based on both relevant and irrelevant documents.
+Unlike their method, we directly model a query-aware irrelevant intent space via dual-space contrastive learning, and enhance the performance with hard negative data augmentation, leading to a more precise intent description.
+2.2 Query-focused Summarization
+In settings where annotated intent descriptions are available, a related task to Q2ID is query-focused summarization (QFS) (Vig et al., 2022; Pagnoni et al., 2023).
+QFS is a subtask of text summarization that aims to generate a summary of one or multiple documents, guided by a query.
+Traditional methods rely on unsupervised extraction, ranking text segments by similarity and query relevance (Wan and Xiao, 2009; Feigenblat et al., 2017).
+Recent QFS datasets (Kulkarni et al., 2020; Fabbri et al., 2022; Zhong et al., 2021) have enabled the rise of QA-driven approaches (Su et al., 2020, 2021).
+More advanced techniques model query relevance through evidence ranking (Xu and Lapata, 2021), latent query optimization (Xu and Lapata, 2022), or pipeline architectures like the coarse-to-fine model in Xu and Lapata (2020).
+To handle long documents, extract-then-abstract strategies (Vig et al., 2022) use sparse attention and segment scoring.
+Other innovations include question-driven pretraining (Pagnoni et al., 2023), contrastive learning (Sotudeh and Goharian, 2023), and joint token-utterance modeling with query-aware attention (Liu et al., 2023a).
+We use QFS models as baselines for the Q2ID task, but there is a fundamental difference between QFS and Q2ID: QFS aims to compress the content of retrieved documents to help users consume information, whereas Q2ID aims to generate a description of what the system likely inferred about the query intent, based on retrieval results.
+We provide a comparison table with related tasks in Appendix A.
+3 Methods
+Figure 2: Overview of our proposed pipeline.
 From left to right, we show (a) Intent-Driven Negative Augmentation method, (b) Contrastive decoder structure with dual cross-attention layers, (c) and (d) Contrastive learning via dual space modeling.
-[…]
-In this section, we present our approach QUIDS that generates descriptive query intents via dual space modeling.
-Figure 2 shows an overview.
-We first introduce the pipeline framework in section 3.1, then the data augmentation method to select intent-driven hard negative samples in section 3.2, and finally the contrastive mechanism via dual space modeling in section 3.3.
-3.1. Pipeline Framework
-We define the contrastive generation task as follows.
-Given a dataset $\mathcal{D}=\{(q,\mathcal{R},\mathcal{I},y)_{j}\}$ with $L$ samples, where $j\in\{0,1,...,L\}$: $q$ is a query, =$\mathcal{R}=\{r_{1},r_{2},...,r_{|\mathcal{R}|}\}$ is a collection of relevant documents for the query, =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ is a collection of irrelevant documents and $y$ is the ground truth query intent (i.e.
-the expected output).
-The goal is to learn the distinctions between relevant and irrelevant inputs based on a query while generating an intent description that exclusively highlights the relevant aspects related to the query.
-To achieve this, our training pipeline consists of 2 steps: (1) Intent-Driven Negative Augmentation (IDNA): mining hard negative documents as irrelevant documents from the entire dataset $D$ based on the query, its relevant document collections, and the ground truth intent, i.e.,
+3.1 Pipeline Framework
+We define the contrastive intent generation task as follows.
+Given a dataset $\mathcal{D}=\{(q,\mathcal{R},\mathcal{I},y)_{j}\}$ with $L$ samples, where $j\in\{0,1,...,L\}$: $q$ is a query, $\mathcal{R}=\{r_{1},r_{2},...,r_{|\mathcal{R}|}\}$ is a collection of relevant documents for the query, $\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ is a collection of irrelevant documents and $y$ is the human-annotated ground truth query intent.
+The modeling goal is to learn the distinctions between relevant and irrelevant inputs based on a query, while generating a system-inferred intent description that exclusively highlights the relevant aspects related to the query.
+To achieve this, our training pipeline consists of 2 steps: (1) Intent-Driven Negative Augmentation (IDNA) and (2) Dual Space Modeling (DualSM).
+3.2 Intent-Driven Negative Augmentation (IDNA)
+The purpose of IDNA is to mine hard negative documents as irrelevant documents from the entire dataset $D$ based on the query, its relevant document collections, and the ground truth intent, i.e.,
 $IDNA(q,\mathcal{R},y,\mathcal{D})=\mathcal{I}^{\prime}$
-where =$\mathcal{I}^{\prime}=\{i^{\prime}_{1},i^{\prime}_{2},...,i^{\prime}_{h}\}$ with $h$ the expected number of irrelevant documents.
-(2) Dual Space Modeling (DSM): contrastively generating a descriptive intent for the query by modeling query-aware relevant and irrelevant intent spaces, i.e.,
-$DSM(q,\mathcal{R},i^{\prime})=\hat{y}$
-3.2. Intent-Driven Negative Augmentation
+where $\mathcal{I}^{\prime}=\{i^{\prime}_{1},i^{\prime}_{2},...,i^{\prime}_{h}\}$ with $h$ the expected number of irrelevant documents.
 Inspired by Liu et al.
@@
 The negative sample augmentation makes the task more challenging for the model, hence improving its discriminative capabilities.
-As shown in Figure 2 (a), we augment hard negative samples using the following steps: For each query $q$ in the training set,
-  1. (1)
-    We rank its relevant documents $\mathcal{R}$ based on their cosine similarity to the concatenated embedding of the query and its ground truth intent $(q;y)$ where (;) denotes concatenation, resulting in $\mathcal{R^{*}}$.
-  2. (2)
-    We concatenate the ranked relevant documents $\mathcal{R^{*}}$ into a single document $R^{*}$ and obtain its embedding $h_{R^{*}}$.
-  3. (3)
-    The original irrelevant documents are initially ranked based on their cosine similarities to the newly concatenated relevant embedding $h_{R^{*}}$.
-  4. (4)
-    If the number of irrelevant documents is smaller than an expected size, the irrelevant document collection $\mathcal{I}^{\prime}$ is augmented by selecting hard negative samples from all other documents in the whole dataset $\mathcal{D}$ according to their cosine similarities to the new embedding $h_{R^{*}}$.
-For step 1, 3 and 4, we use a pre-trained Sentence Transformers model (Reimers and Gurevych, 2019) trained on the MSMARCO Passage Ranking Dataset (Nguyen et al., 2016) to encode the documents.
-In practice, during step 4, we select documents whose cosine similarities exceed a predefined threshold 0.8 until the expected size is reached.
-3.3. Dual Space Modeling
-We use a Transformer-based encoder-decoder architecture, with the BART-large model (Lewis et al., 2020) and the T5-large model (Raffel et al., 2020a)as the backbone, as illustrated in Figure 2 (b).
-To capture the relationship between a query and its relevant and irrelevant documents, we adopt a Siamese dual encoder architecture.
+As shown in Figure 2 (a), we augment hard negative samples for each training query using Algorithm 1.
+Algorithm 1 Intent-Driven Negative Augmentation (IDNA)
+1:query $q$, relevant document collection $\mathcal{R}$, irrelevant document collection $\mathcal{I}$, whole dataset document corpus $\mathcal{D}$, target size $S$, threshold $\tau$
+2:augmented irrelevant documents $\mathcal{I}^{\prime}$ for $q$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+10:while $|\mathcal{I}^{\prime}|<S$ do
+11:  Sample a document $d$ from $\mathcal{D}$
+12:  if $\cos(h_{R^{*}},\text{Encode}(d))>\tau$ then
+13:   Add $d$ to $\mathcal{I}^{\prime}$
+14:  end if
+15:end while
+For encoding, we use a Sentence Transformers model pre-trained on the MSMARCO Passage Ranking dataset (Nguyen et al., 2016).
+In practice, we set 0.8 to the similarity threshold with an analysis of its effect in subsection 5.4.3.
+3.3 Dual Space Modeling (DualSM)
+Dual Space Modeling aims to contrastively generate a descriptive intent for the query by modeling query-aware relevant and irrelevant intent spaces:
+$DualSM(q,\mathcal{R},i^{\prime})=\hat{y}$
+We use a Transformer-based encoder-decoder architecture, with the BART-large model (Lewis et al., 2020) and the T5-large model Raffel et al.
+(2020a) as backbones, as illustrated in Figure 2(b).
+To capture the relationship between a query and its relevant and irrelevant documents, we implement a Siamese dual encoder architecture.
 Based on the encoder outputs, contrastive learning is performed in the representation space to differentiate between embeddings for relevant and irrelevant documents.
 Correspondingly, we design a contrastive decoder to model query-aware relevant and irrelevant intent spaces in a disentangled manner.
-3.3.1. Representation space modeling
-Naturally, we expect the encoder to model and distinguish the relevant documents and irrelevant documents concerning a query in a feature representation space.
-To this end, we implement a dual encoder architecture, with each encoder being a cross-encoder, so that each cross-encoder directly models the relevance of a relevant or irrelevant document given a query by jointly encoding the query-document pair.
-Intuitively, relevant documents are assumed to share similar topics, reflecting their connection to the specific information need behind a query to varying degrees.
-Therefore, it is natural to encode relevant documents collectively by concatenating them according to the ranking outlined in Section 3.2, Step 1.
+3.3.1 Representation Space Modeling (RSM)
+We design a dual cross-encoder architecture to distinguish relevant and irrelevant documents for a given query by jointly encoding each query-document pair.
+Relevant documents, which often share similar topics tied to the query’s intent, are concatenated based on their ranking (Section 3.2, Step 1) and encoded together.
 In contrast, irrelevant documents can be irrelevant to a query in diverse ways, making it impractical to model a meaningful and comprehensive irrelevant feature representation space.
-Instead, we focus on a single irrelevant document $i^{\prime}$ at each training step, using a hard negative sample from the augmented irrelevant document collection $\mathcal{I^{\prime}}$.
-We aim to model a feature representation space for relevant documents, reflecting their shared topics, while ensuring the irrelevant feature representation space is distant from the relevant one due to the dissimilarity of their topics.
-To achieve this, we project both of the encoder last hidden states into a feature representation space via a linear layer.
-The feature representation for each relevant document is obtained by applying average pooling over the token representations corresponding to the document.
-We then pull the relevant feature representations closer together while pushing the irrelevant representation further away from each relevant one (Figure 2 (c)).
-The optimization objective of the relevant feature representation space is to minimize:
-(1) $\mathcal{L}_{rel}=\sum^{k}_{m=1}\sum^{k}_{n=m+1}d(e_{m},e_{n})$
+Therefore, we focus on a single irrelevant document $i^{\prime}$ at each training step, using a hard negative sample from the augmented irrelevant document collection $\mathcal{I^{\prime}}$.
+To model the feature space, we project the encoder’s final hidden states through a linear layer.
+Document embeddings are obtained via average pooling over token representations.
+We optimize the representation space by pulling relevant embeddings closer and pushing the irrelevant one away (Figure 2(c)).
+The objective is to minimize:
+$\mathcal{L}_{rel}=\textstyle\sum^{k}_{m=1}\sum^{k}_{n=m+1}d(e_{m},e_{n})$ (1)
 where $e$ is the embedding of each relevant document, $k$ is the number of relevant documents, and $d$ is a distance function.
@@
 For irrelevant feature representation space, we optimize the margin loss function:
-(2) $\mathcal{L}_{irrel}=\sum^{k}_{m=1}max(t-d(e_{m},\bar{e}),0)$
+$\mathcal{L}_{irrel}=\textstyle\sum^{k}_{m=1}max(t-d(e_{m},\bar{e}),0)$ (2)
 where $\bar{e}$ is the embedding of the irrelevant document, and $t$ is a margin parameter, set to 1 in our case.
 We combine the relevant and irrelevant loss to obtain the encoder contrastive loss as follows:
-(3) $\mathcal{L}_{encoder}=\mathcal{L}_{rel}+\mathcal{L}_{irrel}$
-3.3.2. Disentangling space modeling
-In the previous encoding phase, we obtain contextual feature representations for relevant and irrelevant documents given a query.
-In the decoder part, we want to make sure the decoder generates intent descriptions based on these information sources.
-Hence, we design a contrastive decoder by adding an additional cross attention so that it can attend to both relevant and irrelevant documents.
-To effectively disentangle the contrastive information learned by the model during the decoding process, we apply additional contrastive learning in disentangling space.
-As shown in Figure 2 (d), the model is designed to capture the contrast between relevant and irrelevant information spaces, allowing it to model both relevant and irrelevant intent spaces.
-This enables the generation of more precise and nuanced intent descriptions by focusing on the relevant intent space while minimizing attention on the irrelevant one.
+$\mathcal{L}_{encoder}=\mathcal{L}_{rel}+\mathcal{L}_{irrel}$ (3)
+3.3.2 Disentangling Space Modeling (DSM)
+In decoding, we aim to generate intent descriptions based on the encoded relevant and irrelevant document features.
+To achieve this, we design a contrastive decoder with an added cross-attention layer that attends to both sources.
+To further disentangle relevant from irrelevant information, we apply contrastive learning in a separate disentangling space.
+As shown in Figure 2(d), this helps the model focus on relevant intent while minimizing influence from irrelevant content, enabling more precise and nuanced intent generation.
 Our decoder adopts a Transformer architecture, composed of N identical decoder layers.
@@
 Based on $h_{self,z}^{l}$, we compute relevant document hidden states $h_{rel,z}^{l}$ by applying multi-head attention with cross-attention (MHAtt) to relevant encoder output:
-(4) $h_{rel,z}^{l}=MHAtt(h_{self,z}^{l},h_{R^{*}})$
+$h_{rel,z}^{l}=MHAtt(h_{self,z}^{l},h_{R^{*}})$ (4)
 Similarly, we get the irrelevant document hidden states by attending to irrelevant encoder output:
-(5) $h_{irrel,z}^{l}=MHAtt(h_{self,z}^{l},h_{i^{\prime}})$
+$h_{irrel,z}^{l}=MHAtt(h_{self,z}^{l},h_{i^{\prime}})$ (5)
 From preliminary results, we found that a simple linear combination of $h_{self,z}^{l}$, $h_{rel,z}^{l}$, and $h_{irrel,z}^{l}$ works well to serve as the decoder hidden state to produce the distribution over the target vocabulary:
-(6) $h_{combine,z}^{l}=h_{self,z}^{l}+h_{rel,z}^{l}-h_{irrel,z}^{l}$
-(7) $P^{vocab}_{z}=Softmax(W(h_{combine,z}^{N}))$
-where $W$ indicates a linear transformation. We optimize the model with the negative log likelihood (NLL) objective function to predict the target words:
-(8) $\mathcal{L}_{NLL}=-\sum^{|y|}_{z=1}logP^{vocab}_{z}(y_{z})$
+$h_{combine,z}^{l}=h_{self,z}^{l}+h_{rel,z}^{l}-h_{irrel,z}^{l}$ (6)
+$P^{vocab}_{z}=Softmax(W(h_{combine,z}^{N}))$ (7)
+where $W$ indicates a linear transformation. We optimize the model with the negative log likelihood (NLL) objective to predict the target words:
+$\mathcal{L}_{NLL}=-\textstyle\sum^{|y|}_{z=1}logP^{vocab}_{z}(y_{z})$ (8)
 Corresponding to the representation space contrastive learning, we perform another contrastive learning in the newly proposed disentangling space using hidden states from the last decoder layer.
@@
 We follow the approach of SimCLR (Chen et al., 2020) and use from-batch negative samples $\mathcal{B}$ in the InfoNCE loss (He et al., 2020):
-(9) $\mathcal{L}_{decoder}=-log\frac{exp(cos(f_{c},f_{r})/\tau)}{\sum_{i^{\prime}\in\mathcal{B}}exp(cos(f_{c},f_{i^{\prime}})/\tau)}$
-where $\tau$ is the temperature and $cos(\textperiodcentered,\textperiodcentered)$ defines cosine similarity.
+$\mathcal{L}_{decoder}=-log\frac{exp(cos(f_{c},f_{r})/\tau)}{\sum_{i^{\prime}\in\mathcal{B}}exp(cos(f_{c},f_{i^{\prime}})/\tau)}$ (9)
+where $\tau$ is the temperature and $cos(\cdot,\cdot)$ defines cosine similarity.
 Finally, we combine the original NLL loss together with encoder and decoder loss to obtain the overall loss $\mathcal{L}$ to update all learnable parameters in an end-to-end learning setting:
-(10) $\mathcal{L}_{NLL}=\lambda_{0}\mathcal{L}_{NLL}+\lambda_{1}\mathcal{L}_{Encoder}+\lambda_{2}\mathcal{L}_{Decoder}$
-where $\lambda$ parameters control the balance between the three losses, with their total sum equal to 1.
-4. Experimental Settings
-4.1. Dataset
-Q2ID (Zhang et al., 2020) is a benchmark dataset for query-to-intent description, derived from existing TREC and SemEval collections: TREC is an ongoing series of benchmark activities in IR.
-Q2ID comprises the TREC Dynamic Domain track 2015, 2016, 2017 on dynamic, exploratory search and the TREC 2004 Robust Track on consistency of retrieval technology.
-These tracks together contribute to 510 entries in the Q2ID benchmark dataset.
-SemEval is a series of workshops on Semantic Evaluation.
-Q2ID comprises the English SemEval-2015 Task3 and SemEval-2016 Task3 on Community Question Answering, contributing 4878 entries to the Q2ID dataset.
-For each query in the collections, documents with multi-graded relevance labels were converted into binary labels, indicating whether they are relevant or not to the query.
-In total, the dataset comprises 5,358 entries, each formatted as a quadruple: ¡query, relevant documents, irrelevant documents, intent description¿, with intent descriptions being human-written narratives that provide detailed descriptions of the search intent behind the queries.
-The length statistics are an average of 7.2 tokens per query and 45.5 tokens per intent description.
-We split the 5338 queries into training (5000), validation (100), and test (258) query sets according to the original split setting of the Q2ID benchmark dataset.
-4.2. Baselines
-Since query-focused summarization task and query intent generation task are closely aligned in their focus on user queries and the extraction of relevant information, we evaluate our model against state-of-the-art models from the Q2ID benchmark, QFS task, and additionally include LLM zero-shot and two-shot baselines.
+$\mathcal{L}_{NLL}=\lambda_{0}\mathcal{L}_{NLL}+\lambda_{1}\mathcal{L}_{Encoder}+\lambda_{2}\mathcal{L}_{Decoder}$ (10)
+where the $\lambda$ parameters control the balance between the three losses, with their total sum equal to 1.
+4 Experimental Settings
+4.1 Data
+We conduct experiments on the Q2ID dataset (Zhang et al., 2020), a benchmark for query-to-intent description derived from existing TREC and SemEval collections.
+Specifically, it comprises: TREC: Including the Dynamic Domain tracks (2015–2017) and the 2004 Robust Track, which focus on dynamic, exploratory search and consistency of retrieval technology.
+SemEval: Including the English SemEval-2015 and SemEval-2016 Task 3 tracks on Community Question Answering.
+Q2ID contains a total of 5,358 entries.
+Each entry is structured as a quadruple: <query, relevant documents, irrelevant documents, intent description>, where the intent descriptions are human-written narratives.
+The statistics and more details are provided in Appendix B.1.
+4.2 Baselines
+To reflect the shared focus on user queries and the extraction of relevant content, we compare our model with baselines from four categories: (i) Pretrained Seq2Seq Models: We fine-tune T5-large (Raffel et al., 2020b) and BART-large (Lewis et al., 2020) on the Q2ID dataset.
+BART also serves as the backbone of our QUIDS model.
+(ii) Q2ID Baseline: CtrsGen (Zhang et al., 2020) leverages contrastive generation using a bi-GRU encoder and contrast-weighted attention mechanism.
+(iii) LLM Baseline: We evaluate a instruction-tuned model LLaMA3.1-8B-Instruct (AI, 2024) and a reasoning model OpenAI o3 (OpenAI, 2025) in zero-shot and two-shot settings.
+For the two-shot settings, examples are drawn from TREC and SemEval.
+(iv) QFS Baselines: We include extractive-abstractive models RelReg, RelRegTT (Vig et al., 2022), the segment-based model SegEnc (Vig et al., 2022), the question-driven Socratic (Pagnoni et al., 2023), and the contrast-enhanced Qontsum (Sotudeh and Goharian, 2023).
+Detailed descriptions, training configurations, and reproduction settings for baselines are in Appendix B.4 and B.5.
+4.3 Implementation Details
+Our method is implemented based on the BART-large model (Lewis et al., 2020) using the Huggingface Transformers library (Wolf et al., 2019).
+CrossRel and CrossIrrel attention layers in the decoder are initialized with pre-trained BART weights.
+We optimize the weighted training loss using coefficients $(\lambda_{0}=0.2,\lambda_{1}=0.2,\lambda_{2}=0.6)$ to balance multiple objectives (see Appendix B.2).
+The model is trained with the Adam optimizer, and the final checkpoint is selected based on average ROUGE-{1, 2, L} scores on the validation set.
+We provide additional training details in Appendix B.3.
+4.4 Evaluation Metrics
+We conduct three types of evaluations using different evaluator resources: automatic evaluation, LLM-based evaluation and human evaluation.
+For automatic evaluation, we report recall scores on ROUGE-{1, 2, L} following Zhang et al.
+(2020), along with BERTScore (Zhang et al., 2019), which assesses semantic and syntactic similarity beyond exact word matches.
+We also conduct a human evaluation study using 50 (Sotudeh and Goharian, 2023) randomly selected test samples.
+Five PhD students in Computer Science scored intent descriptions from our model and the best baseline, without knowing which model produced them.
+They rated both models on four customized qualitative criteria with scores ranging from 1 (worst) to 5 (best).
+Four criteria are: (1) Fluency: to what extent the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
+(2) Factual Alignment: to what extent the generated query intent description is factually aligned with the ground truth intent.
+(3) Inclusion score: how well the generated query intent includes important details from the query and relevant documents.
+(4) Exclusion score: how well the generated query intent description excludes information present in the irrelevant documents that is not relevant to the query and relevant documents.
+Inspired by Liu et al.
+(2023b), we also adopt LLM-based evaluation (LLaMa3.1-8B-Instruct and GPT-4o) by prompting instruction-tuned models to assess generations across the same four qualitative metrics.
+Details on the prompt formats, and scoring computation are provided in Appendix F.
+Models RG-1 RG-2 RG-L BS
+[…]
+T5-large 28.87 13.91 23.85 61.64
+BART-large 30.70 13.91 24.63 62.07
+LLaMa3.1 (0) 29.28 7.42 20.90 57.26
+LLaMa3.1 (2) 32.75 9.54 24.34 57.89
+o3 (0) 34.91 6.84 23.93 59.19
+o3 (2) 33.15 6.28 22.81 59.88
+RelReg 26.67 12.83 21.99 59.24
+RelRegTT 27.21 12.77 22.25 59.60
+SegEnc 31.83 14.29 25.18 62.15
++ SOCRATIC Pret. 31.38 13.88 24.91 62.26
+QONTSUM 31.18 14.26 24.87 62.03
+QUIDS_T5 29.40 13.95 24.23 62.00
+QUIDS_BART 34.47 14.86* 26.77* 63.55*
+Table 1: Performance between our model and baselines in terms of automatic evaluation (%).
+$\dagger$ indicates reported performance from previous work.
+‘-’ means the result is inaccessible.
+* indicates the model outperforms the best baseline significantly with paired t-test at $p$-value << 0.05 level.
+Results are averaged over 5 random seeds.
+The best results are highlighted in bold, while the best baseline results are underlined.
+Model RG-1 RG-2 RG-L BS
+QUIDS w/o IDNA 33.48 14.20 25.95 63.17
+QUIDS w/o RSM 34.57 14.39 26.38 63.62
+QUIDS w/o DSM 33.45 13.46 25.88 63.33
+QUIDS 35.95 14.80 27.21 64.33
+Table 2: Ablation study of our QUIDS model with its variants under automatic evaluation (%).
+Figure 3: Distribution of human and LLM evaluation sores on four qualitative metrics.
+5 Experimental Results
+5.1 Overall Results
+We compare model performance between QUIDS and baselines in Table 1.
+The results show that: (1) QUIDS outperforms all baselines except o3 on RG-1, including those with larger model sizes, indicating that the gains stem primarily from our dual-space modeling design rather than model size or pretraining knowledge; (2) Our approach is compatible with both T5 and BART architectures.
+Notably, BART-large outperforms T5-large despite having nearly half the model size; (3) The QFS models that we implemented for the query intent generation task outperform the Q2ID-specific baseline, CtrsGen.
+QUIDS further significantly outperforms the best QFS model, SegEnc; (4) The two-shot setting with the LLaMa3.1-8B-Instruct model significantly outperforms the zero-shot setting in ROUGE scores, while showing only minor improvements in BERTScore.
+This suggests that without fine-tuning, generated intents may be lexically similar but semantically misaligned; (5) While o3 shows strong reasoning and generates richer descriptions, it does not outperform our model on ROUGE and BERTScore.
+We observed that o3 often paraphrases intent with different wording, leading to lower ROUGE scores.
+Its longer outputs and inclusion of detailed justifications or summaries may dilute the concise intent signal.
+This suggests that strong reasoning alone may not align well with the goal of generating concise, system-inferred intent.
+5.2 Ablation Study
+We perform an ablation study based on the BART-Large model to evaluate the contribution of key components in our approach under three settings: (1) without the IDNA module (w/o IDNA), (2) without contrastive learning in the encoder (w/o RSM), and (3) without contrastive learning in the decoder (w/o DSM).
+Results are shown in Table 2.
+Excluding contrastive learning from the decoder leads to the largest performance drop, underscoring its role in modeling a discriminative intent space.
+Removing it from the encoder results in a smaller decline, suggesting that representation space modeling still contributes to relevance awareness.
+Figure 4: Boxplot of human and LLM evaluation scores on 4 metrics of our model on different intent types.
+5.3 Human and LLM Evaluation
+We assess the quality of generated intents from QUIDS and the best baseline SegEnc using both human and LLM-based evaluations.
+Inter-annotator agreement, measured by weighted Cohen’s $\kappa$, and LLM-human correlations, measured by Spearman and Kendall $\tau$, indicate fair to moderate consistency across metrics (Table 7, Appendix C).
+As shown in Table 3, QUIDS outperforms SegEnc on all metrics except Factual Alignment, where humans prefer SegEnc.
+Further analysis (subsection 5.4.1) shows this stems from the dominance of informational queries, on which SegEnc performs better.
+QUIDS, by contrast, performs better on exploratory queries.
+Figure 3 further illustrates score distributions, revealing three key insights:
+Method Model Fluen. Align. Inclu. Exclu.
+Human SegEnc 4.75 3.90 3.94 4.77
+QUIDS 4.80 3.80 4.06 4.80
+GPT-4o SegEnc 3.70 2.64 2.67 3.83
+QUIDS 3.71 2.79 2.69 4.00
+LLaMa3.1 SegEnc 3.25 2.91 3.17 4.07
+QUIDS 3.48 3.17 3.42 4.11
+Table 3: Comparison of human evaluation and LLM evalaution in terms of Fluency, Factual Alignment, Inclusion score and Exclusion score.
+(1) Human scores are generally higher than LLM scores, especially for Fluency and Exclusion.
+The larger variability for fluency scores suggests humans may tolerate minor fluency issues.
+(2) Human evaluations show broader and lower score distributions for Factual Alignment and Inclusion, aligning more with LLaMa3.1 (Table 7).
+In contrast, they mirror GPT-4o’s narrower distribution for Fluency and Exclusion, where correlation is higher.
+This suggests that evaluators differ in how they assess each metric.
+(3) Fluency and Factual Alignment show stronger alignment between LLM and human evaluations, likely due to being less context-dependent.
+In contrast, Inclusion and Exclusion scores exhibit weaker correlations, indicating inconsistencies in evaluating context-sensitive criteria.
+Intent RG-1 RG-2 RG-L BS
+Informational 35.69 14.51 26.82 63.88
+Exploratory 41.55 23.24 38.28 76.65
+Table 4: Comparison of automatic evaluation on our model for different intent types.
+Figure 5: Case study indicating the role of contrastive examples in the decoder stage.
+Token-level decoder cross-attention weights are shown for a generated intent token (red) are shown with (a) and without (b) an irrelevant document in the model input.
+Deeper color indicates a higher value.
+5.4 In-depth Analysis
+5.4.1 Analysis of Intent Types
+We classify the queries according to their underlying search intent into two categories: (1) Informational Intent: Natural language questions seeking detailed information or solutions, typically longer and contextual.
+Queries from the SemEval dataset fall under this category.
+(2) Exploratory Intent: Term-based queries aimed at broad exploration with minimal context or structure.
+Queries from the TREC datasets are categorized here.
+Our automatic and human evaluation by intent type indicates that QUIDS is indeed more successful for exploratory tasks than for informational intent tasks.
+Results are shown in Table 4 and Figure 4.
+See Appendix D for detailed analysis.
+Figure 6: Robustness analysis on (a) input length and (b) IDNA threshold.
+5.4.2 Analysis of Document Length
+In real-world scenarios, relevant documents associated with a query can vary in length.
+Documents over 1024 tokens are truncated in our architecture.
+To assess robustness to input length, we group test documents into short (<512 tokens), medium (512–1024), and long (>1024).
+In the test dataset, 204 are short, 25 medium, and 29 long.
+Figure 6(a) shows that the overall differences in performance among all the input length categories are minimal across all metrics, suggesting that 1024-token inputs are sufficient for inferring the underlying query intent.
+This may explain why long-document QFS models offer limited gains in this task.
+5.4.3 Sensitivity Analysis of IDNA threshold
+Intent-driven negative augmentation (IDNA) selects irrelevant documents with high similarity to relevant ones as hard negatives for contrastive learning.
+We noticed that most document similarity scores fall between 0.6 and 0.9, making the method robust even with lower thresholds (e.g., <0.6).
+Performance improves as the threshold increases within this range, with 0.8 yielding the best results, which we use in our experiments.
+5.5 Case Study
+Figure 5 illustrates how the model uses cross-attention in the decoder stage to identify irrelevant semantics from a low-ranked document.
+When generating ‘impacted’ without an irrelevant document (Figure 5(a)), the model focuses on economic effects on cars, indicated by ‘price’ in Intent 1.
+With an irrelevant document in Figure 5(b), while similar economic attentions are observed across both relevant and irrelevant documents when generating the word ‘effects’ in Intent 2, the model successfully identifies tokens related to prices and cars in relevant documents as irrelevant.
+This demonstrates the model’s ability to filter out irrelevant content using contrastive learning in the decoder.
+We include another failure case study in Appendix E.
+6 Conclusions
+We introduced a novel dual-space modeling approach for the query intent generation task.
+Our approach implements contrastive learning in both encoding and decoding phases, combined with intent-driven hard negative augmentation during data preprocessing, to automatically generate detailed and precise intent descriptions, surfacing what the system likely inferred the query to mean.
+Experimental results show that our model can effectively filter out irrelevant information from the relevant intent space, leading to more accurate intent descriptions than all baselines, including models for Query-Focused Summarization.
+In future work, we plan to improve contextual understanding in distinguishing relevant from irrelevant information and extend our approach to conversational search by mining exploratory needs and explaining the understanding of query intents.
+Our long-term aim is to improve the transparency in the retrieval process, in particular for exploratory search needs.
+Acknowledgments
+This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
+Limitations
+Training Efficiency Trade-offs.
+Augmenting irrelevant documents enhances robustness but linearly increases training time.
+We therefore limited negative documents to three per query, which partially alleviates this trade-off but remains suboptimal.
+To fundamentally resolve this efficiency bottleneck, two promising directions are: (1) adaptive dynamic sampling that prioritizes high-impact negatives through real-time gradient analysis, and (2) curriculum-based augmentation progressively introducing harder negatives as training stabilizes.
+Dataset Imbalance.
+Informational queries dominate the training data over exploratory ones.
+While our model shows promising performance in exploratory search scenarios, this bias limits deeper intent analysis.
+Future work should expand out experiments to more datasets, focusing on exploratory queries.
+One option would be to use LLM-generated synthetic data, specifically creating pseudo-documents that mimic multi-faceted exploratory intents.
+This approach maintains intent modeling consistency while enabling systematic investigation of query complexity, without requiring manual annotation efforts.
+Appendix A Detailed Comparison with Related Work
+A detailed comparison between our task—Query intent description generation —and related tasks such as Query Understanding (QU), Query-Focused Summarization (QFS), and Pseudo-Relevance Feedback (PRF) is provided in Table 6.
+Appendix B Experimental Setting Details
+B.1 Dataset Details
+In constructing the Q2ID dataset, documents with multi-graded relevance labels were converted into binary labels, indicating whether each document is relevant to the query. The dataset is composed of:
+  • 510 entries from TREC tracks (Dynamic Domain 2015–2017, Robust 2004)
+  • 4,878 entries from SemEval-2015/2016 Task 3 on Community Question Answering
+Each data point is formatted as a quadruple: <query, relevant documents, irrelevant documents, intent description>.
+The average query length is 7.2 tokens, and the average intent description length is 45.5 tokens.
+We follow the original split of Q2ID: 5,000 queries for training, 100 for validation, and 258 for testing.
+B.2 Sensitivity Analysis of Loss Weights
+We conduct a sensitivity analysis through a grid search over different combinations of the loss weights ($\lambda_{0}$, $\lambda_{1}$, $\lambda_{2}$).
+The results of several representative configurations are summarized in Table 5.
+$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ RG-1 RG-2 RG-3 BS
+- - - 34.53 14.42 26.85 63.33
+0.33 0.33 0.33 33.36 14.64 26.22 63.45
+0.20 0.20 0.60 34.51 15.36 27.13 63.52
+0.30 0.30 0.40 34.17 14.31 26.30 63.41
+0.10 0.10 0.80 32.37 14.59 25.81 62.87
+0.20 0.40 0.40 33.20 14.56 26.10 63.22
+Table 5: Sensitivity analysis of loss weights. ‘-’ indicates that the $\lambda$ parameters were treated as learnable during training.
+We observed that configurations assigning relatively higher weight to $\lambda_{2}$ tend to yield better overall performance.
+Overly large $\lambda_{2}$ values (e.g., 0.8) degrade performance by reducing the contribution of the other loss components.
+Configurations that assign more weight to $\lambda_{0}$ or $\lambda_{1}$ alone lead to inferior performance, sometimes resulting in no valid generation, and are therefore omitted for clarity.
+Based on these observations, we adopt $(\lambda_{0},\lambda_{1},\lambda_{2})=(0.2,0.2,0.6)$ as the default setting.
+Task Query Understanding (Classification / Clustering / Expansion) Query-Focused Summarization (QFS) Pseudo-Relevance Feedback (PRF) Our Task: Query Intent Generation (QIG)
+Goal Predict query intent classes, discover latent topics, or expand queries for better retrieval performance.
+Summarize relevant documents to help users consume content.
+Refine or reformulate queries to improve retrieval performance.
+Generate a natural language description of the search system’s inferred intent behind a query.
+Output Form Labels (e.g., informational/navigational), clusters, or expanded query terms.
+Natural language summary (abstractive or extractive).
+Modified query or re-weighted terms.
+Natural language explanation of inferred query intent.
+Use of Irrelevant Documents Not used.
+Focus is on query-only or top-ranked documents.
+Rarely used; mainly uses pseudo-relevant documents.
+Not used; PRF assumes top-ranked documents are relevant.
+Explicitly contrasts relevant and irrelevant documents for intent disentanglement.
+Application Stage Pre-retrieval; typically before document scoring.
+Post-retrieval summarization.
+Interleaved or pre-retrieval (used for re-ranking or expansion).
+Post-retrieval; supporting user query refinement and retrieval debugging
+User Utility Improves ranking accuracy and personalization; not visible to users.
+Helps users consume content more efficiently.
+Improves recall or relevance through backend query rewriting.
+Helps users understand potential mismatches between their intended query meaning and the system’s inferred intent.
+Table 6: Comparison between Query Understanding (QU) tasks, Query-Focused Summarization (QFS), Pseudo-Relevance Feedback (PRF), and our Query Intent Generation (QIG) task.
+B.3 Implementation
+To balance efficiency and effectiveness during IDNA augmentation, we set the expected number of irrelevant documents per query to three for training.
+This results in augmenting 1,984 queries with at least three irrelevant documents each.
+Training is conducted for 10 epochs using the Adam optimizer with a learning rate of 0.0001.
+During decoding, we set a maximum sequence length of 256 tokens and apply beam search with a beam size of 4.
+We also set a no-repeat n-gram size of 3 to reduce redundancy.
+B.4 Baselines
 Pretrained sequence-to-sequence model baselines: T5 (Raffel et al., 2020b): a Transformer-based encoder-decoder (Vaswani et al., 2017) model trained on a diverse and extensive dataset.
@@
 During decoding, it computes contrast scores by considering irrelevant documents to adjust sentence-level attention weights in the relevant documents.
-  • Large Language Model (LLM) baseline: LLama3.1 (AI, 2024): We use the Llama3.1-8B instruction-tuned text-only model in both zero-shot and two-shot settings and conduct five experimental runs for each setting.
+Large Language Model (LLM) baseline: LLaMa3.1-8B-Instruct (AI, 2024): We evaluate the LLaMa3.1-8B instruction-tuned text-only model under both zero-shot and two-shot settings and conduct five experimental runs for each setting.
 For two-shot setting, we randomly using two different examples per run – one sourced from TREC and the other from SemEval.
+OpenAI o3 (OpenAI, 2025): We evaluate the reasoning-focused OpenAI o3 model.
+While the model internally generates reasoning tokens, we focus our analysis on the visible output tokens.
+The same experimental settings and number of runs are used for comparison.
 Query Focused Summarization (QFS) baselines: RelReg (Vig et al., 2022) and RelRegTT (Vig et al., 2022): two-step approaches for QFS consisting of an score-and-rank extractor and an abstractor.
@@
 We train the QFS baselines on the Q2ID dataset using the original code provided by the authors, except for Qontsum, which we independently reproduced.
-4.3. Implementation Details
-We build our method based on the BART-large (Lewis et al., 2020) model using Huggingface Transformers (Wolf et al., 2019).
-CrossRel attention and CrossIrrel attention layers in the decoder blocks share the same initial weights from the pre-trained BART-large model.
-For the IDNA step, we set the expected number of irrelevant documents per query to three in the training set.
-To ensure this, we augment 1,984 queries so that each query includes at least three irrelevant documents.
-We train the model for 10 epochs using the Adam optimizer with a learning rate of 0.0001.
-The final checkpoints are chosen based on the average ROUGE-{1, 2, L} scores against the ground truth query intent obtained from the validation set.
-During decoding, we set the maximum length to 256 tokens and apply beam search with a beam size of 4, using a no-repeat n-gram size of 3.
-We optimize $\lambda$ during training and fix on $(\lambda_{0}=0.2,\lambda_{1}=0.2,\lambda_{2}=0.6)$ to balance the three losses.
-All experiments are conducted five times, and we report the average results across these runs.
-The implementation details for the other baseline models can be found in Appendix A.1.
-4.4. Evaluation Metrics
-We conduct three kinds of evaluations using different evaluator resources: automatic evaluation, LLM-based evaluation and human evaluation.
-For automatic evaluation, we report recall scores on ROUGE-{1, 2, L} following Zhang et al.
-(2020), along with BERTScore (Zhang et al., 2019), which assesses semantic and syntactic similarity beyond exact word matches.
-Additionally, we conduct a human evaluation study using 50 randomly selected test samples.
-Three PhD students in Computer Science scored intent descriptions from our model and the best baseline, without knowing which model produced them.
-They rated both models on four customized qualitative criteria, with scores ranging from 1 (worst) to 5 (best), and the final scores were averaged across the three annotators.
-  • Fluency: to what extent the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
-  • Factual Alignment: to what extent the generated query intent description is factually aligned with the ground truth intent.
-  • Inclusion score: how well the generated query intent includes important details from the query and relevant documents.
-  • Exclusion score: how well the generated query intent description excludes information present in the irrelevant documents that is not relevant to the query and relevant documents.
-However, conventional automatic evaluation metrics tend to show weak correlation with human judgments.
-Recent study (Fu et al., 2024; Liu et al., 2023b) propose using LLMs as evaluators for natural language generation (NLG) systems, scoring generated outputs based on generation probability rather than reference targets.
-Following the method of (Liu et al., 2023b), we use LLaMa3.1-8B [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ as instruction-tuned evaluators to assess the generated intent across four qualitative metrics.
-Specifically, we define the evaluation task and criteria, prompting the LLM to generate chain-of-thoughts (CoT) for the ‘Evaluation Steps’.
-For LLaMa3.1-8B, we use the output token probabilities from the LLMs to normalize the scores and take their weighted summation as the final results: $score=\sum_{i=1}^{n}p(s_{i})\times s_{i}$ where $S=\{s_{1},s_{2},...,s_{n}\}$ represents the predefined score set from the prompt, with a maximum value of 5 in our case.
-For the close-sourced GPT-4o, we sample 20 times to estimate the token probabilities.
-Example prompts are in Appendix B.
-5. Experimental Results
-Table 1.
-Performance between our model and baselines in terms of automatic evaluation (%).
-$\dagger$ indicates reported performance from previous work.
-‘-’ means the result is inaccessible.
-* indicates the model outperforms the best baseline significantly with paired t-test at $p$-value $<$ 0.05 level.
-The best results are highlighted in bold, while the best baseline results are underlined.
-Models RG-1 RG-2 RG-L BS
-[…]
-T5-large 28.87 13.91 23.85 61.64
-BART-large 30.70 13.91 24.63 62.07
-LLaMa3.1 zero-shot 29.28 7.42 20.90 57.26
-LLaMa3.1 2-shot 32.75 9.54 24.34 57.89
-RelReg 26.67 12.83 21.99 59.24
-RelRegTT 27.21 12.77 22.25 59.60
-SegEnc 31.83 14.29 25.18 62.15
-+ SOCRATIC Pret. 31.38 13.88 24.91 62.26
-QONTSUM 31.18 14.26 24.87 62.03
-QUIDS_T5 29.40 13.95 24.23 62.00
-QUIDS_BART 34.47* 14.86* 26.77* 63.55*
-5.1. Overall Results
-We compare model performance between QUIDS and baselines in Table  1.
-The results show that: (1) Our model QUIDS outperforms all other baselines; (2) Our approach is compatible with both T5 and BART architectures, which both have a transformer-based encoder-decoder structure.
-This demonstrates the effectiveness of our design in the query intent generation task.
-Notably, BART-large outperforms T5-large despite having nearly half the model size, highlighting its efficiency in this task.
-(3) We implemented QFS models for the query intent generation task, which already outperform the baseline model, CtrsGen.
-However, our design further improves ROUGE-1 by 2.64 and ROUGE-L by 1.59 compared to the best QFS model, SegEnc; (4) In our experiments, the two-shot setting with the LLaMa3.1 8B model significantly outperforms the zero-shot setting in ROUGE scores, while showing only minor improvements in BERTScore.
-This suggests that without fine-tuning, intent generation may be lexically similar but semantically misaligned.
-5.2. Ablation Study
-Table 2. Ablation study of our QUIDS model with its variants under automatic evaluation (%).
-Models RG-1 RG-2 RG-L BS
-QUIDS w/o IDNA 33.48 14.20 25.95 63.17
-QUIDS w/o RSM 34.57 14.39 26.38 63.62
-QUIDS w/o DSM 33.45 13.46 25.88 63.33
-QUIDS 35.95 14.80 27.21 64.33
-We conduct an ablation study on the test set to assess the impact of various modules in our proposed approach under three conditions: (1) We remove the IDNA module to examine the effect of data preprocessing (w/o IDNA); (2) We remove contrastive learning from the encoder and (3) decoder separately to assess the importance of representation space modeling (w/o RSM) and disentangling space modeling (w/o DSM), respectively.
-Table 2 shows the results.
-Removing IDNA leads to a 1.16-point drop in BERTScore, suggesting that incorporating harder irrelevant documents helps the model capture relevant intent information more effectively.
-Removing contrastive learning from the decoder leads to a significant drop in all metrics, highlighting its importance in modeling a distinguishable intent space and generating accurate intent descriptions.
-Finally, removing contrastive learning from the encoder results in a slight decrease, indicating that representation space modeling aids the model in distinguishing relevant from irrelevant information.
-Due to the similar architecture and experimental performance, we omit the T5 model results.
-5.3. Human and LLM Evaluation
-We analyze the quality of the generated intents by our model QUIDS and the best baseline model SegEnc with human evaluation and LLM-based approaches.
-Human inter-annotator agreement is assessed with weighted Cohen’s kappa, revealing fair to moderate consistency across the four metrics.
-As shown in table 3, QUIDS consistently receives higher scores from all three evaluators, except for Factual Alignment, where human evaluators prefer SegEnc’s outputs.
-Table 3. Comparison of human evaluation and LLM evalaution in terms of Fluency (Fluen.), Factual Alignment (Align.), Inclusion score (Inclu.) and Exclusion score (Exclus.).
-Method Model Fluen. Align. Inclu. Exluc.
-Human SegEnc 4.71 3.70 3.88 4.71
-QUIDS 4.73 3.61 3.97 4.73
-GPT-4o SegEnc 3.70 2.64 2.67 3.83
-QUIDS 3.71 2.79 2.69 4.00
-LLaMa3.1 SegEnc 3.25 2.91 3.17 4.07
-QUIDS 3.48 3.17 3.42 4.11
-Figure 3. Distribution of human and LLM evaluation sores on four qualitative metrics.
-[…]
-Figure 4.
-(a) Visualizes the token-level decoder cross-attention weights for a relevant document snippet when generating the token ‘impacted’ in Intent 1, without an irrelevant document provided.
-(b) for relevant document and irrelevant document when generating the token ‘effects’ in Intent 2, with an irrelevant document provided.
-Deeper color indicates a higher value.
-[…]
-To examine the performance differences between humans and LLMs in this task, we present boxplots in Figure 3 that display the score distributions for each metric.
-Additionally, we calculate the Spearman and Kendall-Tau correlations between the evaluations from LLMs and human judgments, as shown in Table 5 in Appendix A.2.
-From this analysis, we derive the following insights:
-  • Humans tend to assign higher scores than GPT-4o and LLaMa 3.1 across all metrics, particularly for Fluency and Exclusion, where scores nearly reach 5.
-The variability in fluency scores suggests humans are more tolerant of less fluent text.
-  • Human evaluations exhibit the narrowest range, while GPT-4o shows a similar range but with more critical scores.
-As supported by the correlation table in Appendix A.2, GPT-4o aligns more closely with human evaluations, particularly in Factual Alignment.
-However, a difference in model preferences for this metric prompts further analysis of sub-dataset differences in Section 5.4.
-  • For the metrics Factual Alignment and Inclusion, all methods tend to exhibit a wider distribution with generally lower scores.
-This suggests variability in the assessments of these two metrics, potentially indicating inconsistency in how both human and LLM evaluators interpret the scoring criteria.
-5.4. Analysis of Intent Types
-For analysis of the effect of intent type, we classify the queries according to their underlying search intent into two categories: (1) Informational Intent: Natural language questions seeking detailed information or solutions, typically longer and contextual.
-Queries from the SemEval dataset fall under this category.
-(2) Exploratory Intent: Term-based queries aimed at broad exploration with minimal context or structure.
-Queries from the TREC datasets are categorized here.
+B.5 Implementation Details for Baselines
+RelReg and RelRegTT share the same abstractor, a BART-large model, which also serves as the backbone model for SegEnc, Socratic, and Qontsum.
+For RelReg and RelRegTT, we use an input segment length of 1024, whereas SegEnc-based models utilize an input segment length of 512, with a total input length of 4096.
+For Socratic training, we use the checkpoint pretrained on Books3 (Csaky and Recski, 2020) from the Huggingface Model Hub [^2] https://huggingface.co/Salesforce/socratic-books-30M and fine-tune it on Q2ID dataset using SegEnc mechanism.
+We reproduce the work of Qontsum with the segment length of 512 tokens, temperature of 0.6 and ($\lambda_{0}=0.6,\lambda_{1}=0.2,\lambda_{2}=0.2$) in joint learning.
+For all models that divide input text into segments, we apply a 50% overlap between each segment and its adjacent one.
+Correlation Model Fluency Alignment Inclusion Exclusion Average
+$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
+Corr (Human, GPT-4o) SegEnc 0.320 0.252 0.387 0.295 0.158 0.121 0.375 0.309 0.310 0.244
+QUIDS 0.476 0.375 0.495 0.386 0.265 0.199 0.361 0.286 0.399 0.312
+Corr (Human, LLaMa3.1) SegEnc 0.224 0.173 0.434 0.329 0.153 0.116 0.343 0.272 0.289 0.223
+QUIDS 0.373 0.294 0.557 0.424 0.365 0.262 0.158 0.126 0.363 0.276
+Table 7: Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations between Human evaluation and LLM evaluation of different metrics.
+Method Intent Fluen. Align. Inclu. Exclu.
+Human Info. 4.78 3.71 4.04 4.80
+Expl. 4.89 4.19 4.14 4.81
+GPT-4o Info. 3.64 2.70 2.63 3.95
+Expl. 4.02 3.17 2.96 4.20
+LLaMa3.1 Info. 3.33 2.96 3.35 3.99
+Expl. 3.72 4.11 3.72 4.64
+Table 8: Comparison of Human and LLM evaluation on informational and exploratory intent types on our model.
+Appendix C Correlation with Human Evaluation
+We assess the correlation between human and LLM evaluators across four qualitative evaluation criteria, presenting the Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations for the best SOTA model SegEnc and our QUIDS model in Table 7.
+Overall, our model demonstrates significantly higher human correspondence across all metrics compared to SegEnc, with the exception of the Exclusion score.
+Correlation performance varies by metric; for Fluency and Factual Alignment—criteria requiring less contextual information—there is a relatively higher degree of agreement with human evaluations.
+In contrast, the Inclusion and Exclusion scores, which depend on diverse and contextual sources, show lower correlation, suggesting that humans and LLM evaluators adopt different evaluation strategies for more complex criteria.
+Additionally, we observe that different LLM evaluators exhibit human-like evaluation behaviors across various metrics: LLaMa3.1 shows greater human correspondence in Factual Alignment and Inclusion scores, whereas GPT-4o aligns more closely with human evaluations in Fluency and Exclusion scores.
+Appendix D Evaluation on Intent Types
+In Table 3, we observe a human preference over the SegEnc model on metric Factual Alignment, which measures how well the generated query intent description is factually aligned with the ground truth intent.
+We guess it is due to the model performance difference on different sub-datasets, or on different intent types.
+And hence we further analyse the evaluations on different intent types.
+Automatic Evaluation
 In the 258 test samples, there are 20 queries with exploratory intents and 238 with informational intents.
@@
 This is an aspect that makes our approach more suitable for exploratory search rather than informative search.
-We additionally conduct a breakdown analysis by intent types for the human and LLM results in Appendix A.3.
-Table 4. Comparison of automatic evaluation on our model QUIDS for different intent types.
-Intent RG-1 RG-2 RG-L BS
-Informational 35.69 14.51 26.82 63.88
-Exploratory 41.55 23.24 38.28 76.65
-5.5. Analysis of Document Length
-In real-world scenarios, relevant documents associated with a query can vary in length.
-To assess the impact of document length on the generated output, we categorize the documents into three groups based on the BART-large input limit (1024 tokens): documents shorter than 512 tokens are classified as short documents; those ranging from 512 to 1024 tokens are termed medium-sized documents; and documents longer than 1024 tokens, which the model will truncate, are labeled as long documents.
-In our test dataset, there are 204 short, 25 medium, and 29 long documents.
-The overall differences in performance among all the input length categories are minimal across all metrics (also see Figure 6 in Appendix A.4).
-This indicates that 1024 tokens provide sufficient information for our model to effectively capture the underlying query intent.
-This may explain why QFS approaches, particularly those designed to handle lengthy documents, do not demonstrate significant advantages in the query intent generation task.
-5.6. Case Study
-To intuitively understand how irrelevant documents influence our model, we compare the generated intents for a query ‘Freon-12’ from test set with and without providing an irrelevant document.
-Figure 4 illustrates the token-level decoder cross-attention weights for both a relevant document snippet, and an irrelevant document snippet when provided.
-As shown in Figure 4 (a), without an irrelevant document provided, when generating the word ‘impacted’, the model’s focus is centered on the economic effects on cars, as indicated by the mention of ‘price’ in Intent 1.
-With an irrelevant document provided in Figure 4 (b), when generating the word ‘effects’ in Intent 2, similar attentions can be detected on ‘prices’, ‘\$’, ‘costs’, ‘auto’ for both relevant and irrelevant documents.
-However, as shown in Generated Intent 2, our model now treats tokens related to prices and cars in relevant documents as irrelevant.
-This demonstrates that out model effectively generates intent descriptions that excludes this unrelated information.
-6. Conclusion
-In this paper, we introduced a novel dual-space modeling approach for the task of query intent generation.
-Our approach implements contrastive learning in both the encoding and decoding phases, combined with intent-driven hard negative augmentation during data preprocessing, to automatically generate detailed and precise intent descriptions.
-Experimental results show that our model effectively filters out irrelevant information from the relevant intent space, leading to more accurate intent descriptions and greater transparency in the retrieval process.
-For future work, we aim to extend our approach to conversational search by first mining exploratory needs and then explaining the understanding of query intents.
-Acknowledgements.
-This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
-Appendix A Appendix A
-A.1. Implementation details for baseline models
-RelReg and RelRegTT share the same abstractor, a BART-large model, which also serves as the backbone model for SegEnc, Socratic, and Qontsum.
-For RelReg and RelRegTT, we use an input segment length of 1024, whereas SegEnc-based models utilize an input segment length of 512, with a total input length of 4096.
-For Socratic training, we use the checkpoint pretrained on Books3 (Csaky and Recski, 2020) from the Huggingface Model Hub [^3] https://huggingface.co/Salesforce/socratic-books-30M and fine-tune it on Q2ID dataset using SegEnc mechanism.
-We reproduce the work of Qontsum with the segment length of 512 tokens, temperature of 0.6 and ($\lambda_{0}=0.6,\lambda_{1}=0.2,\lambda_{2}=0.2$) in joint learning.
-For all models that divide input text into segments, we apply a 50% overlap between each segment and its adjacent one.
-A.2. Correlation with Human Evaluation
-Table 5. Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations between Human evaluation and LLM evaluation of different metrics.
-Correlation Model Fluen. Align. Inclu. Exluc.
-$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
-Corr (Human, GPT-4o) SegEnc 0.377 0.304 0.409 0.310 0.184 0.159 0.369 0.306
-QUIDS 0.476 0.375 0.517 0.412 0.252 0.199 0.361 0.286
-Corr (Human, LLaMa3.1) SegEnc 0.243 0.187 0.459 0.344 0.143 0.111 0.224 0.177
-QUIDS 0.343 0.278 0.553 0.423 0.325 0.248 0.158 0.126
-We assess the correlation between human and LLM evaluators across four qualitative evaluation criteria, presenting the Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations for the best SOTA model SegEnc and our QUIDS model in Table 5.
-Overall, our model demonstrates significantly higher human correspondence across all metrics compared to SegEnc, with the exception of the Exclusion score.
-Correlation performance varies by metric; for Fluency and Factual Alignment—criteria requiring less contextual information—there is a relatively higher degree of agreement with human evaluations.
-In contrast, the Inclusion and Exclusion scores, which depend on diverse and contextual sources, show lower correlation, suggesting that humans and LLM evaluators adopt different evaluation strategies for more complex criteria.
-Additionally, we observe that different LLM evaluators exhibit human-like evaluation behaviors across various metrics: LLaMa3.1 shows greater human correspondence in Factual Alignment and Inclusion scores, whereas GPT-4o aligns more closely with human evaluations in Fluency and Exclusion scores.
-A.3. Breakdown Analysis on Intent Types
-Table 6. Comparision of Human and LLM evaluation on informational and exploratory intent types on our model.
-Method Intent Fluen. Align. Inclu. Exluc.
-Human Info. 4.71 3.51 3.98 4.73
-Expl. 4.85 4.04 3.88 4.74
-GPT-4o Info. 3.64 2.70 2.63 3.95
-Expl. 4.02 3.17 2.96 4.20
-LLaMa3.1 Info. 3.33 2.96 3.35 3.99
-Expl. 3.72 4.11 3.72 4.64
-Figure 5. Boxplot of evaluation scores on 4 metrics of our model on different intent types.
-[…]
-In Table 3, we observe a human preference over the SegEnc model on metric Factual Alignment, which measures how well the generated query intent description is factually aligned with the ground truth intent.
-We guess it is due to the model performance difference on different sub-datasets, or on different intent types.
-And hence we further analyse the human evaluations on different intent types.
-Table 6 present human and LLM evaluations on our model regarding two intent types.
-While humans prefer our model for exploratory intent with 4.04 (QUIDS) vs.
-3.81 (SegEnc), SegEnc is favored for informative intent with 3.51 (QUIDS) vs.
+Human and LLM Evaluation
+Table 8 presents human and LLM evaluations on our model regarding two intent types.
+In general, exploratory intents consistently outperform informational intents across all metrics.
+This finding, derived from 50 test samples, aligns with automatic evaluation results on the full test dataset (Table 4).
+Figure 4 shows the evaluation score distribution by intent type on our model, compared to the overall model performance in Figure 3.
+The informational intent distribution closely mirrors the overall performance, suggesting that informational queries dominate the dataset and largely influence performance.
+However, exploratory queries, despite being less frequent, demonstrate superior performance in this task.
+When diving into the factual alignment in Figure 3, while humans prefer our model for exploratory intent with 4.19 (QUIDS) vs.
+3.94 (SegEnc), SegEnc is favored for informative intent with 3.71 (QUIDS) vs.
 3.89 (SegEnc).
-Since informative intent queries dominate, this leads to a lower average score for our model.
-We also present the evaluation scores distribution on different intent types over 50 the same test samples in subsection 5.4.
-Figure 5 is a breakdown of Figure 3 on our model for four metrics.
-Based on Table 6 and Figure 5, we have following insights:
-  • Exploratory intent types generally outperform informational intent types across all metrics, except for human evaluations on inclusion.
-This finding, based on 50 test samples from both human and LLM evaluations, is consistent with the automatic evaluation results for the full test dataset, as shown in Table 4.
-  • The boxplots for informational intent on the right side of each metric in Figure 5 exhibit a distribution similar to the overall performance of our model presented in Figure 3, suggesting that informational intent queries dominate the dataset and largely drive performance outcomes.
-However, exploratory intent queries outperform in this particular task, indicating their stronger performance despite being less frequent.
-A.4. Breakdown Analysis by document length
-Figure 6 shows the results of the automatic evaluation, broken down based on document length.
-Figure 6. Comparison of automatic evaluation on our model QUIDS on different input lengths.
-[…]
-Appendix B Appendix B
-B.1. Evaluation prompt on Fluency
+Since informative intent queries dominate, this leads to a lower average score for our model on this metric.
+These findings indicate that our model is well-suited for exploratory search.
+Appendix E Failure Case Study
+Figure 7: Failure case study.
+Token-level decoder cross-attention weights for a generated intent token (red) are shown with (c) and without (d) an irrelevant document.
+Deeper color indicates a higher value.
+Figure 7 illustrates a failure example of filtering irrelevant information when an irrelevant document is provided.
+The token-level decoder cross-attention weights are compared when generating a content word in the intent, with (c) and without (d) an irrelevant document.
+When generating the keyword ‘UK’ and ‘Dubai’, the model mainly focuses on ‘petrol’ and ‘cigarettes’ in the relevant documents for both (c) and (d), which are also contextually important in the generated intent.
+However, the model fails to recognize the relationship between ‘middle east’ in the irrelevance document and ‘Dubai’, leading to the unwanted inclusion of ‘Dubai’ in the intent 2.
+This highlights that our model may struggle with excluding information that requires commonsense reasoning or domain-specific knowledge.
+A direction for future work is to develop advanced approaches that enhance contextual understanding for complex scenarios.
+Appendix F LLM-based Evaluation Details
+Following the method of (Liu et al., 2023b), we use LLaMa3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct.
+and GPT-4o [^4] https://openai.com/index/hello-gpt-4o/.
+as instruction-tuned evaluators to assess the generated intent across four qualitative metrics.
+Specifically, we define the evaluation task and criteria, prompting the LLM to generate chain-of-thoughts (CoT) for the ‘Evaluation Steps’.
+For LLaMa3.1-8B-Instruct, we use the output token probabilities from the LLMs to normalize the scores and take their weighted summation as the final results:
+$score=\sum_{i=1}^{n}p(s_{i})\times s_{i}$ (11)
+where $S=\{s_{1},s_{2},...,s_{n}\}$ represents the predefined score set from the prompt, with a maximum value of 5 in our case.
+For the close-sourced GPT-4o, we sample 20 times to estimate the token probabilities.
+An example prompt for each model is presented below.
+F.1 General Evaluation Prompt
 You will be given a query, relevant and irrelevant documents with respect to the query.
@@
 Evaluation Criteria:
-Fluency Score (1-5) - This metric measures if the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
+<<MetricCriteria>>
 Evaluation steps:
-1. Carefully review the provided query, relevant, and irrelevant documents to understand the context and content.
-2. Read the ground truth query intent description to understand the ideal response. This serves as a benchmark for evaluating the fluency of the generated description.
-3. Carefully read the generated query intent description. Focus on the fluency aspect, considering factors such as grammatical correctness, naturalness, clarity, coherence, and readability.
-Assign a rating from 1 to 5 based on the level of fluency.
+<<EvaluationSteps>>
 Query:
@@
 Evaluation Form (scores ONLY):
+- <<MetricName>>:
+F.2 Evaluation Prompt on Fluency
+Evaluation Criteria:
+Fluency Score (1-5) - This metric measures if the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
+Evaluation steps:
+1. Carefully review the provided query, relevant, and irrelevant documents to understand the context and content.
+2. Read the ground truth query intent description to understand the ideal response. This serves as a benchmark for evaluating the fluency of the generated description.
+3. Carefully read the generated query intent description. Focus on the fluency aspect, considering factors such as grammatical correctness, naturalness, clarity, coherence, and readability.
+Assign a rating from 1 to 5 based on the level of fluency.
+Evaluation Form (scores ONLY):
 - Fluency:
-B.2. Evaluation prompt on Factual Alignment
-You will be given a query, relevant and irrelevant documents with respect to the query.
-You will also be given a generated query intent description based on the query and documents.
-The ground truth query intent description will also be provided.
-Your task is to rate the query intent description on one metric.
-Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
+F.3 Evaluation Prompt on Factual Alignment
 Evaluation Criteria:
@@
 Assign a rating from 1 to 5 based on the level of factual alignment.
-Query:
-{{Query}}
-Relevant documents:
-{{Relevant documents}}
-Irrelevant documents:
-{{Irrelevant documents}}
-Generated Intent:
-{{Generated intent}}
-Ground Truth Intent:
-{{Gound truth intent}}
 Evaluation Form (scores ONLY):
 - Factual Alignment:
-B.3. Evaluation prompt on Inclusion Score
-You will be given a query, relevant and irrelevant documents with respect to the query.
-You will also be given a generated query intent description based on the query and documents.
-The ground truth query intent description will also be provided.
-Your task is to rate the query intent description on one metric.
-Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
+F.4 Evaluation Prompt on Inclusion Score
 Evaluation Criteria:
@@
 Assign a rating from 1 to 5 based on the thoroughness and relevance of the included details.
-Query:
-{{Query}}
-Relevant documents:
-{{Relevant documents}}
-Irrelevant documents:
-{{Irrelevant documents}}
-Generated Intent:
-{{Generated intent}}
-Ground Truth Intent:
-{{Gound truth intent}}
 Evaluation Form (scores ONLY):
 - Inclusion Score:
-B.4. Evaluation prompt on Exclusion Score
-You will be given a query, relevant and irrelevant documents with respect to the query.
-You will also be given a generated query intent description based on the query and documents.
-The ground truth query intent description will also be provided.
-Your task is to rate the query intent description on one metric.
-Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
+F.5 Evaluation Prompt on Exclusion Score
 Evaluation Criteria:
@@
 Assign a rating from 1 to 5 based on the level of exclusion of irrelevant details.
-Query:
-{{Query}}
-Relevant documents:
-{{Relevant documents}}
-Irrelevant documents:
-{{Irrelevant documents}}
-Generated Intent:
-{{Generated intent}}
-Ground Truth Intent:
-{{Gound truth intent}}
 Evaluation Form (scores ONLY):